### v1

# Pedagogical Alignment of Large Language Models

Abstract
In this paper, we introduce the novel concept of pedagogically aligned Large Language Models (LLMs) that signifies a transformative shift in the application of LLMs within educational contexts. Rather than providing direct responses to user queries, pedagogically-aligned LLMs function as scaffolding tools, breaking complex problems into manageable subproblems and guiding students towards the final answer through constructive feedback and hints. The objective is to equip learners with problem-solving strategies that deepen their understanding and internalization of the subject matter. Previous research in this field has primarily applied the supervised finetuning approach without framing the objective as an alignment problem, hence not employing reinforcement learning through human feedback (RLHF) methods. This study reinterprets the narrative by viewing the task through the lens of alignment and demonstrates how RLHF methods emerge naturally as a superior alternative for aligning LLM behaviour. Building on this perspective, we propose a novel approach for constructing a reward dataset specifically designed for the pedagogical alignment of LLMs. We apply three state-of-the-art RLHF algorithms and find that they outperform SFT significantly. Our qualitative analyses across model differences and hyperparameter sensitivity further validate the superiority of RLHF over SFT. Also, our study sheds light on the potential of online feedback for enhancing the performance of pedagogically-aligned LLMs, thus providing valuable insights for the advancement of these models in educational settings.
1 Introduction
Large Language Models (LLMs) are an integral part of the modern digital landscape, and ensuring their alignment to be honest, harmless, and helpful - the ‘3H’ principle - is critical to their effective usage [1]. However, the needs of a student learning a subject or preparing for an exam differ substantially from those of a casual internet user. While the latter might seek immediate answers to their queries, the former benefits more from a structured, guided approach that encourages active learning [2] and critical thinking [3]. Therefore, if we aim to leverage LLMs to assist students’ learning, we must redefine the ‘helpfulness’ criterion in a way that specifically supports this demographic [4].
Figure 1: The image depicts the comparison between a traditional Large Language Model (LLM) interaction (left) and a pedagogically-aligned LLM interaction (right). The traditional LLM directly provides the user with the answer, while the pedagogically-aligned LLM guides the student to the solution by presenting a series of subproblems. This elucidates the concept of pedagogical alignment, emphasizing the transformation from direct problem-solving to a guided, scaffolded learning experience.
In this paper, we propose a novel concept of pedagogical alignment for LLMs which aligns LLMs to function as scaffolding tools [5], breaking complex problems into manageable subproblems and guiding students towards the final answer through constructive feedback and hint [6]. This approach redefines the ‘helpfulness’ criterion of the ‘3H’ principle, tailoring it to the unique needs and learning objectives of students. We argue that to truly assist students, LLMs should not merely provide immediate answers, but rather guide learners through a structured, step-by-step process that encourages active learning and critical thinking. The pedagogical alignment of LLMs is not just a theoretical construct; it is grounded in the belief that learning can be significantly improved when students are given the opportunity to actively participate in their learning process [7] guided by feedback [8].
Our work builds upon the framework of Conversational Learning with Analytical Step-by-Step Strategies (CLASS) [9, 10], a tool that applies above principles of learning sciences to construct an Intelligent Tutoring System (ITS) using LLMs. However, their approach is implemented through supervised fine-tuning of the LLM, which we identify as a limitation. We argue that the desired behavior to foster learning can be more effectively modeled using reinforcement learning. However, the challenge lies in the absence of a reward dataset necessary for reinforcement learning in education context.
To address this challenge, we reinterpret the CLASS framework from a pedagogical alignment perspective, proposing that it can be viewed as a potential reinforcement learning problem where the rewards are associated with pedagogical alignment. This perspective allows us to construct a preference (reward) dataset, laying the groundwork to enhance the alignment of LLMs to better serve educational purposes.
Next, using this reward dataset, we study the effective of three RLHF algorithms: Direct Preference Optimization (DPO) [11], Identity Preference Optimization (IPO) [12], and Kahneman Tversky Optimization (KTO) [13] for pedagogical alignment of two state-of-the-art open-source LLMs - Mistral [14] and Zephyr [15]. We measure the the efficacy of these algorithms on pedagogical alignment by an accuracy metric. This metric quantifies the alignment by considering whether the correct pedagogical action was taken by the model - a crucial aspect of our pedagogical alignment approach.
Our findings reveal that DPO and KTO significantly enhance the performance of both Mistral and Zephyr models, improving our pedagogical alignment accuracy by 37 and 23 points respectively, compared to CLASS’ SFT approach. Also, we examine the influence of the beta hyperparameter on the performance of these alignment algorithms since beta is a critical parameter that significantly affects the outcome of the alignment process.
Our contributions in this paper are threefold:
  1. 1.
    We introduce the concept of “pedagogical alignment,” a new approach to align LLMs in a manner that fosters optimal student learning.
  2. 2.
    We reinterpret the CLASS framework from a pedagogical alignment perspective, which allows us to create a reward dataset.
  3. 3.
    Using the reward dataset, we demonstrate the effectiveness of reinforcement learning-based alignment algorithms on state-of-the-art LLMs.
Through this work, we aim to contribute to the effective use of LLMs in education, highlighting their potential as productive learning aids and enhancing the learning experience for students.
Table 1: Inference prompt of CLASS framework.
Prompt: Instructions to act as a Tutorbot, an AI-powered chatbot designed to help student with a question. Your goal is to break problem into  subproblems for the student, and walk the student to the final answer using the subproblems and hints.
For each student response, think which category your response falls in, and then frame the reply using this json structure:
{ "Evaluation": "a,b,c,d,e,f,g",
   "Action Based on Evaluation": "1,2,3,4,5,6,..,12",
   "Subproblem State": "x,y,z",
   "Subproblem": "..",
   "Tutorbot": "..", }
Evaluation of Student Response:
a) Correct Response
b) Incorrect Response
c) Partially Correct Response
d) Ambiguous or Unclear Response
e) ...
Tutorbot Actions:
If "a" is the evaluation, then:
1) Confirm the correct answer. Check for completeness of the answer. If solution is incomplete, notify the student to complete the solution.
Prompt continues: If "b" is the evaluation, then: 2) Notify the student about the mistake, Provide constructive feedback to pinpoint the errors, Offer hints 3) Step in to provide a solution if the student is unable to answer even after multiple attempts. If "c" is the evaluation, then: 4) Acknowledge the accurate parts, Promptly notify the student about the mistake, Provide constructive feedback to pinpoint the errors, Offer helpful hints 5) Step in to provide a solution if the student is unable to answer after multiple attempts If "d" is the evaluation, then: 6) Actively seek clarification through relevant follow-up questions. Request student to provide specific information Subproblem State: x) One of the subproblems is currently being solved y) Subproblem finished, move to next subproblem z) Subproblem finished, no next subproblem, problem finished
2 CLASS through the lens of pedagogical alignment
This section presents the concept of pedagogical alignment and uses it to examine the CLASS framework. This examination highlights the potential of application of reinforcement learning and paves the way for constructing a reward dataset to enhance the pedagogical alignment of LLMs.
2.1 Pedagogical alignment
In the landscape of LLMs, the term “alignment" is a concept used to describe the process of adapting LLMs to align with human preferences. In the context of online conversational agents, alignment specifically refers to making an LLM honest, harmless, and helpful to a typical internet user [1]. However, the needs of a student are distinct from those of a regular internet user. As illustrated in Figure 1, when an internet user poses a question to a LLM, the LLM’s response should directly address the user’s query. This is the primary use case for out-of-the-box LLMs like Mistral [14] and Llama [16], which are designed for such interactions. However, to promote optimal student learning, LLMs should guide the student to reach the correct answer through their own reasoning by employing effective scaffolding [17], involving problem decomposition, constructive feedback, and timely hints.
Here, we introduce our unique perspective of viewing the CLASS framework in the context of pedagogical alignment. The CLASS tutorbot differentiates itself from conventional out-of-the-box LLMs through its approach to student queries. Instead of directly providing answers, the CLASS tutorbot takes on a more interactive and guided approach. It begins by breaking the problem into smaller, manageable subproblems. Each subproblem presents a learning opportunity for the student, guiding them towards the final solution in a step-by-step manner.
The CLASS framework achieves effective scaffolding through a defined discrete action and state space. For every student response, the CLASS tutorbot determines whether the student’s response is correct, incorrect, or ambiguous. For instance, if the response is incorrect, an out-of-the-box LLM’s instinctual response would be to provide the correct answer to the user. However, the CLASS tutorbot’s initial reaction is to offer constructive feedback to the student and allow the student to arrive at the right answer independently. If the student is unable to do so, then the CLASS’s eventual response is to step in and provide the correct answer (refer to action 2 and 3 in inference prompt  1).
Similarly, if the student’s response is partially correct, the CLASS tutorbot’s first reaction is to inform the student about the correct aspects and then highlight the errors, encouraging the student to correct their answer independently (action 4 in prompt 1), as opposed to providing the complete answer, which is a typical reaction of an out-of-the-box LLM.
Pedagogical alignment: We define the term ‘pedagogical alignment’ to describe this shift in behavior that CLASS introduces, in contrast to an out-of-the-box LLM, to promote learning [18]. While out-of-the-box LLMs are designed to be helpful to general users, pedagogically aligned LLMs are tailored to foster optimal learning experiences for students.
2.2 Reinforcement learning algorithms for alignment
While the CLASS framework used SFT for pedagogical alignment, recent developments in the field of Reinforcement Learning from Human Feedback (RLHF) present promising alternatives. RLHF refines LLM’s behavior based on human-derived preferences or feedback, promoting a more profound congruence with human ethical standards and objectives. The seminal work by [19] introduced the concept of leveraging human preferences to steer the fine-tuning process of language models, thereby laying the foundational stone for RLHF. Subsequent advancements [20, 21] have led to notable enhancements in the RLHF pipeline, augmenting alignment efficiency and overall model performance.
The RLHF methodology unfolds in a structured three-step process:
  1. 1.
    Supervised Fine-Tuning (SFT): Initially, the LLM undergoes SFT using a task-specific dataset. This step adjusts the model’s initial parameters to align with the downstream task’s objectives, providing a tailored foundation for further refinement.
  2. 2.
    Preference Modeling: Following SFT, the model generates multiple responses under varied generation parameters. Human evaluators then rank these responses based on their preferences. This ranking process informs the development of a preference model, which learns to assign scores to LLM outputs, effectively translating human judgments into quantifiable metrics.
  3. 3.
    Online Reinforcement Learning: Utilizing the preference model, LLMs are further aligned through online reinforcement learning algorithms such as Proximal Policy Optimization (PPO) [22]. This phase iteratively refines the model’s parameters to maximize the alignment with human feedback, enhancing the model’s performance and ethical alignment.
More recently, Direct Preference Optimization (DPO) [11] has emerged as a streamlined and robust advancement over RLHF, offering superior performance by forgoing the necessity for a reward model. DPO enhances alignment efficiency by directly optimizing a preference-based loss function, thereby simplifying the implementation and the operational efficiency. DPO is designed to directly leverage a dataset of preferences, represented as ${(x,y_{w},y_{l})}$, where each tuple consists of a prompt $x$, a preferred response $y_{w}$, and a dis-preferred response $y_{l}$. Let, $\pi_{\theta}$ be the LLM being finetuned, and $\pi_{\text{ref}}$ be a reference LLM (generally an SFTed model). Then, DPO’s optimization problem can be written as:
$L_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)\right],$ (1)
where $\sigma$ is the sigmoid function, and $\beta$ is an algorithmic parameter. However, DPO has been observed to over-fit to the preference dataset, and is believed to be sensitive to hyperparameter adjustments. To prevent the over-fitting behaviour of DPO’s optimization algorithm, recent proposals include Identity Preference Optimisation (IPO) [12] and Kahneman-Tversky Optimization (KTO) [13].
All the three methodologies, DPO, IPO, and KTO, require a dataset composed of preferences, denoted as ${(x,y_{w},y_{l})}$. The assembly of such datasets in a real-world context is notably labor-intensive and financially burdensome. However, in the subsequent section, we introduce our innovative reward/preference data generation approach, which effectively circumvents these challenges, paving the way for efficient and cost-effective pedagogical alignment of LLMs.
3 Preference Data Generation for Pedagogical Alignment
Figure 2: This illustration shows the process of generating pedagogically-aligned preference data using the CLASS framework. The GPT-student asks a question, to which both the GPT-tutor and SFT-tutor respond. The key focus here is on the divergence in the ‘Action Based on Evaluation’ between the two tutors. In this example, the GPT-tutor’s response is deemed more pedagogically aligned because it encourages the student to engage in critical thinking and attempt the problem again, instead of directly providing the correct answer. This action mismatch between the tutor responses allows us to construct a preference dataset that distinguishes between the pedagogically preferred (chosen) and less effective (rejected) responses. The preference dataset serves as the foundation for applying reinforcement learning-based alignment algorithms, with the goal of enhancing the pedagogical alignment of large language models.
In this section, we describe our DPO, IPO, and KTO compatible data generation pipeline. The objective is to generate a preference dataset $\mathcal{D}_{p}$ which necessitates distinguishing between pedagogically chosen and rejected tutor responses based on their alignment with desired scaffolding strategies. $\mathcal{D}_{p}$ will mimic a dataset structure that is compatible with DPO, IPO, and KTO to optimally represent these preferences. This dataset has three fundamental components: context, chosen, and rejected. In our case, the context represents the accumulated conversational history between the student and the tutor leading up to a particular interaction point, providing the necessary backdrop for the tutor’s next dialogue turn. The chosen contains the pedagogically preferred or optimal tutor responses to the context, while rejected includes those tutor responses considered less effective or misaligned with the scaffolding approach.
CLASS Conversational Dataset $\mathcal{D}_{c}$ and Agents: Our base dataset is the CLASS conversational dataset represented by $\mathcal{D}_{c}=\{(\mathbf{x}_{i},\mathbf{y}_{i})|i=1,\ldots,N\}$, comprising $N$ student-tutor conversational turns. Each $\mathbf{x}$ represents a sequence of student utterances, and each corresponding $\mathbf{y}$ represents the tutor response.
To generate the preference dataset $\mathcal{D}_{p}$ from $\mathcal{D}_{c}$, we employ three autonomous agents: GPT-student ($\mathcal{A}_{\text{G}}$), GPT-tutor ($\mathcal{A}_{\text{T}}$), and SFT-tutor ($\mathcal{A}_{\text{S}}$). The conversation starts with GPT-student $\mathcal{A}_{\text{G}}$ posing a question ($Q$) sourced from a question bank to which both GPT-tutor $\mathcal{A}_{\text{T}}$ and SFT-tutor $\mathcal{A}_{\text{S}}$ provide responses in the following JSON structure:
$\mathbf{y}_{i}=\left\{\begin{array}[]{l}\text{``Evaluation of Student Response": ``a, b, c, \ldots",}\\ \text{``Action Based on Evaluation": ``1, 2, 3, \ldots, 12",}\\ \text{``Subproblem State": ``x, y, z",}\\ \text{``Subproblem": ``...",}\\ \text{``Tutorbot": ``..." }\end{array}\right.$ (2)
Defining Context: We define the context $C_{t}$ for any given response at time $t$ as all preceding conversational turns up to that point:
$C_{t}=\{(\mathbf{x}_{i}^{(t)},\mathbf{y}_{i}^{(t-1)})|i<t\},$ (3)
where $\mathbf{x}_{i}^{(t)}$ denotes the sequence of student response till time $t$, and $\mathbf{y}_{i}^{(t-1)}$ denotes all tutor responses up to, but not including, the response at time $t$. This construction provides a detailed backdrop for each next tutor interaction, feeding into the decision mechanism for chosen and rejected responses.
Preference Dataset $\mathcal{D}_{p}$ Construction: Given this understanding of context, $\mathcal{D}_{p}$ is structured to encapsulate interactions where the pedagogically preferred (chosen) response diverges from the alternative (rejected) based on specific pedagogical criteria. This divergence is identified by a function $f$ that assesses the pivotal attributes of each response. Hence, $\mathcal{D}_{p}$ is defined as:
[…]
$\text{and }f(\mathbf{y}_{t}^{\mathcal{A}_{T}})\neq f(\mathbf{y}_{t}^{\mathcal{A}_{S}}),\forall t\in\{1,\ldots,N\}\big{\}}$ (4)
where $C_{t}=x,y_{w}=y_{choosen},y_{l}=y_{rejected}$ from DPO equation (1) from section 2.2, and the function $f$ is the key to distinguishing between chosen and rejected tutor responses. $f$ indicates a discrepancy between $\mathcal{A}_{\text{T}}$’s and $\mathcal{A}_{\text{S}}$’s responses if any of the fields “Evaluation", “Action Based on Evaluation", or “Subproblem State" diverge between $\mathcal{A}_{\text{T}}$ and $\mathcal{A}_{\text{S}}$, $f$ asserts $\mathcal{A}_{\text{T}}$’s response as preferable, under the assumption that $\mathcal{A}_{\text{T}}$ offers a more accurate pedagogical model, alongside the rejected response from $\mathcal{A}_{\text{S}}$. This condition is captured through the function $f$—mapping responses to their pedagogical attributes:
$f(\mathbf{y}_{i}^{\mathcal{A}})=\left(\mathrm{Eval}(\mathbf{y}_{i}^{\mathcal{A}}),\,\mathrm{ActionEval}(\mathbf{y}_{i}^{\mathcal{A}}),\,\mathrm{SubprobState}(\mathbf{y}_{i}^{\mathcal{A}})\right)$ (5)
where,
  • $\text{Eval}(\mathbf{y}_{i}^{\mathcal{A}})$ extracts insights on the tutor’s grading of student’s solution or question understanding.
  • $\text{ActionEval}(\mathbf{y}_{i}^{\mathcal{A}})$ outlines the tutor’s recommended actions based on the evaluation.
  • $\text{SubprobState}(\mathbf{y}_{i}^{\mathcal{A}})$ provides an understanding of the tutor’s perception of the student’s current understanding or progress on sub-problems within a larger problem context.
The evaluation of these attributes by $f$ facilitates an objective basis to deem one response as chosen (preferable) and another as rejected (less aligned with desired pedagogical outcomes) through a systematic assessment of their pedagogical value and alignment.
4 Experiments
In this section, we provide a comprehensive overview of our dataset construction, experimental design, and the subsequent findings derived from the application of state-of-the-art RLHF algorithms to train pedagogically-aligned LLMs.
4.1 Dataset and Evaluation
Our experimental design requires an extensive conversational dataset to train and test SFT and RLHF models effectively. The CLASS framework [9] used GPT-4 to generate 610 conversations to train their SFT model. However, we need to generate a larger volume of conversations to train RLHF models and conduct comprehensive testing. Following the strategy outlined by CLASS, we set about creating these additional conversations. We used the CLASS scaffolding prompts to generate the problems, subproblems, and associated hints, which were based on biology topics from the OpenStax college book [23], consistent with CLASS’ context. These problems seed the context of the conversations generated through GPT-4 in the next step using the CLASS conversational prompt. In these simulated student-tutor conversations, the tutor uses the subproblems and hints to guide the student toward the final answer. Through this methodology, we successfully generated an additional 1128 conversations, resulting in 1738 conversations. Each conversation, on average, comprises approximately 8.2 rounds and 1311.0 words. This comprehensive dataset provides a solid platform for training and comparing SFT and RLHF models.
To create the training and testing set for SFT and RLHF, we randomly sampled our conversational dataset into 3 partitions: the SFT training dataset $D_{c}$ (600 conversations), the RLHF training dataset $D_{p}$ (600 conversations), and the test dataset $D_{t}$ (450 conversations). As each round within the conversations is processed into a separate data point, $D_{c}$ contains 4942 data points, $D_{p}$ contains 4921 data points, and $D_{t}$ contains 3701 data points.
For the evaluation, we focus on the three classification fields within the model JSON responses described in section 3: Evaluation of Student Response, Action Based on Evaluation, and Subproblem State. We employ accuracy in these three fields to measure the pedagogical alignment of LLMs since these metrics offer a comprehensive view of the model’s performance, evaluating the correctness of the responses, the appropriateness of the actions, and the state of problem-solving. Hence, LLMs that achieve higher accuracy in these fields are considered to provide more pedagogically aligned student assistance.
Table 2: SFT vs DPO, IPO, KTO. Each cell shows the average accuracy across three classification fields - Evaluation of student response, Action based on the evaluation, and Subproblem State, respectively. We observed that for both Mistral (Mistral-7B-Instruct-v0.2) and Zephyr (zephyr-7b-beta), all three RL alignment algorithms - DPO, KTO, and IPO - showed notable accuracy gains in accuracy across all fields as compared to SFT. For both models, DPO and KTO outperformed IPO. IPO performed best for Vicuna (vicuna-7b-v1.5) compared to the other two.
Model SFT DPO IPO KTO
Mistral 35.0 (34,34,37) 74.0 (72,72,78) 57.7 (56,56,61) 74.7 (73,73,80)
Vicuna 74.3 (73,70,80) 75.0 (73,72,80) 76.0 (74,72,82) 75.3 (73,72,81)
Zephyr 10.7 (12,11,9) 74.0 (72,71,79) 21.7 (22,22,21) 74.7 (73,72,79)
Table 3: LLM Performance on key benchmarks. This table presents the performance of Mistral (Mistral-7B-Instruct-v0.2), Zephyr (zephyr-7b-beta), and Vicuna(vicuna-7b-v1.5) on different NLP benchmarks [24, 25]. We selected Mistral and Zephyr for our experiments based on their superior performance across various benchmarks, which provided a thorough and fair assessment across a spectrum of reasoning and general knowledge tasks. We chose Vicuna because it served as the base model for the CLASS framework.
Model Average ARC HellaSwag MMLU TruthfulQA Winogrande GSM8K
Mistral 65.71 63.14 84.88 60.78 68.26 77.19 40.03
Zephyr 61.95 62.03 84.36 61.07 57.45 77.74 29.04
Vicuna 52.06 53.24 77.39 50.82 50.33 72.06 8.11
4.2 Models and and Training Details
We experimented on 3 different LLMs: Mistral (Mistral-7B-Instruct-v0.2), Zephyr (zephyr-7b-beta), and Vicuna (vicuna-7b-v1.5). We selected Mistral and Zephyr based on the their top performance on diverse NLP benchmarks (Table 3). Among all, several benchmark in particular evaluate scientific knowledge and education-based question answering capabilities: ARC (multiple choice questions from grade school science exams) [26], MMLU (multiple choice questions from 57 high school and college subjects across STEM, the humanities, the social sciences, and more) [27], and GSM8K (linguistically diverse grade school math word problems) [28]. These benchmarks offer a comprehensive, unbiased evaluation on the models’ performance, especially in education setting. Furthermore, despite Vicuna’s inferior performance compared to Mistral and Zephyr, we included it in our selection because it was the base model for the CLASS framework’s initial experiments and represented the state of the art at its launch.
Our experiments included training and evaluating SFT and RLHF models. It took around 22 NVIDIA A6000 GPU hours for one cycle (SFT + Evaluation + RLHF + Evaluation). A total of 17 cycles, which are 374 GPU hours, are needed to complete all the experiments. We conducted the SFT experiments using FastChat [29]. For the choice of SFT hyperparameters, we refered to parameters used to instruct fine-tune Vicuna [29]: We employed a learning rate of 2e-5 with the AdamW optimizer [30], a batch size of 16, a cosine scheduler, a weight decay of 0.05, a warmup ratio of 0.1, and a total number of 3 epochs. We conducted the RLHF experiments using TRL [31]. For the choice of RLHF hyperparamters, we referred to parameters used to align Zephyr [32]: We employed a learning rate of 1e-7 with the AdamW optimizer, a batch size of 16, a cosine scheduler, a weight decay of 0.05, a warmup ratio of 0.1, and a total number of 3 epochs.
4.3 Experiment 1: SFT vs DPO, IPO, and KTO
In this experiment, we compared three alignment algorithms: Direct Preference Optimization (DPO), Identity Preference Optimization (IPO), and Kahneman Tversky Optimization (KTO) on the SFT models. We measured the accuracy across three classification fields in the responses: Evaluation of Student Response, Action based on Evaluation, and Subproblem State.
We used the same RLHF training setting as described in Section 4.2 with a beta of 0.1 for DPO, IPO, and KTO across all models. After applying the alignment algorithms, we observed significant accuracy improvements on Mistral and Zephyr SFT models. Specifically, Mistral demonstrated improvements of 39%, 22.7%, and 39.7% from DPO, IPO, and KTO respectively, while Zephyr showed improvements of 63.3%, 11%, and 64% respectively (Table 2). Notably, for both Mistral and Zephyr, DPO and KTO significantly outperformed IPO. However, the Vicuna model, already noted for its high SFT performance, showed only marginal improvements from the alignment; all three algorithms exhibited roughly similar performance. Our results are in alignment with the findings of similar studies [32], which have reported superior performance of DPO and KTO over IPO on Zephyr when the model was fine-tuned on preference datasets, such as Intel’s ORCA dataset pairs [33] and the Ultra-Feedback dataset [15].
This experiment underscored the pivotal role of alignment algorithms in improving the accuracy of SFT models. Mistral and Zephyr, the top performing models, exhibited pronounced improvements over Vicuna. Therefore, we demonstrated the efficacy of pedagogical alignment on state of the art models.
4.4 Experiment 2: Effect of Beta on DPO and IPO
Figure 3: DPO and IPO performance as a function of beta. This figure depicts how the beta hyper-parameter affects the average classification accuracy of DPO and IPO aligned models. We found that both DPO and IPO performances were significantly influenced by beta adjustments. Specifically, a beta of 0.3 optimized results for DPO, with Mistral and Zephyr achieving around 75% accuracy, while a higher beta of 0.9 notably reduced their accuracy. In contrast, IPO performed best with a beta of 0.1, and increasing the beta decreased the accuracy.
In this experiment, we analyzed the impact of the beta parameter on the performance of DPO and IPO alignment algorithms. Beta is a key hyperparameter in all three alignment algorithms (DPO, IPO, and KTO). It controls the strength of the Kullback-Leibler (KL) divergence penalty between the trained model and the reference model within the loss function. Nevertheless, the choice of beta depends on the specific model and dataset and thus necessitates an empirical study.
Similar to experiments in Section 4.3, we measured the average accuracy across the three classification fields. We found that the performance of DPO and IPO is sensitive to the choice of beta. As depicted in figure 3, for DPO, we observed that a beta value of 0.3 yielded the best results: Mistral and Zephyr achieved an average accuracy of 74.7% and 75%. Conversely, a larger beta value, such as 0.9, diminished the performance of the models. Zephyr suffered a large loss of 25.7 % in average accuracy from using 0.9 instead of 0.3. As opposed to IPO, we observed that a beta of 0.1 produced the best performance (figure 3). A larger choice of beta in IPO almost always came with a loss of performance. As beta increases, Mistral’s accuracy dropped from 57.7% to 36 %, while Zephyr’s accuracy dropped from 21.7 % to 11.3 %. Our results corroborate the evidence from related studies [32], demonstrating that various alignment algorithms exhibit distinct sensitivities to beta values, necessitating tailored beta selections.
5 Conclusion
In this paper, we have introduced and explored the concept of pedagogical alignment of LLMs to align these models with educational objectives that are aimed at fostering optimal student learning. By reinterpreting the CLASS framework as a reinforcement learning problem, we constructed a reward dataset, thereby providing a foundation for enhancing the pedagogical alignment of LLMs. Our experimental results, derived from applying reinforcement learning-based alignment algorithms on state-of-the-art open-source LLMs, corroborate our hypothesis. The RLHF methods we employed have proven to outperform the SFT approach significantly, favoring the alignment of LLMs with pedagogical objectives. Furthermore, our study highlights the potential of online feedback for enhancing the performance of pedagogically-aligned LLMs. In conclusion, while our research has made significant strides in aligning LLMs with educational needs, further exploration and refinement of reinforcement learning methods for this purpose remain necessary. We anticipate that our work will serve as a stepping stone towards the optimal integration of LLMs in educational settings, paving the way for future research in this exciting domain.
Acknowledgments
This work was supported by NSF grants 1842378, ONR grant N0014-20-1-2534, AFOSR grant FA9550-22-1-0060, and a Vannevar Bush Faculty Fellowship, ONR grant N00014-18-1-2047.

### v2

@@
 Abstract
-In this paper, we introduce the novel concept of pedagogically aligned Large Language Models (LLMs) that signifies a transformative shift in the application of LLMs within educational contexts.
-Rather than providing direct responses to user queries, pedagogically-aligned LLMs function as scaffolding tools, breaking complex problems into manageable subproblems and guiding students towards the final answer through constructive feedback and hints.
-The objective is to equip learners with problem-solving strategies that deepen their understanding and internalization of the subject matter.
-Previous research in this field has primarily applied the supervised finetuning approach without framing the objective as an alignment problem, hence not employing reinforcement learning through human feedback (RLHF) methods.
-This study reinterprets the narrative by viewing the task through the lens of alignment and demonstrates how RLHF methods emerge naturally as a superior alternative for aligning LLM behaviour.
-Building on this perspective, we propose a novel approach for constructing a reward dataset specifically designed for the pedagogical alignment of LLMs.
-We apply three state-of-the-art RLHF algorithms and find that they outperform SFT significantly.
-Our qualitative analyses across model differences and hyperparameter sensitivity further validate the superiority of RLHF over SFT.
-Also, our study sheds light on the potential of online feedback for enhancing the performance of pedagogically-aligned LLMs, thus providing valuable insights for the advancement of these models in educational settings.
+Large Language Models (LLMs) have shown remarkable potential in various domains, including education.
+However, their application in educational contexts requires careful ‘pedagogical alignment’, where these models are tuned to assist students in a step-by-step problem-solving process, rather than providing direct answers.
+In this paper, we investigate algorithms for learning from human preferences (LHP) to achieve this alignment objective.
+A key challenge in this process is the lack of preference datasets to guide the alignment.
+To address this, we propose a novel approach for constructing such a dataset using synthetic data generation techniques.
+Leveraging this dataset, our experiments reveal the superiority of LHP methods over standard supervised fine-tuning (SFT) for pedagogical alignment.
+Applying LHP algorithms to Mistral and Zephyr, we observe a significant improvement in their ability to guide students, enhancing the pedagogical alignment accuracy by 37.0 and 23.0 points, respectively, compared to SFT.
+Additionally, we introduce a unique approach for evaluating the pedagogical effectiveness of LLMs using perplexity analysis.
+We curate a second preference dataset, comparing the perplexity of LLMs’ responses when offering scaffolded step-by-step guidance versus direct solutions.
+This allows us to quantify the extent to which LLMs prioritize pedagogically effective actions in their generation.
+Our analysis provides compelling evidence for the superiority of LHP methods over SFT in optimizing LLMs’ behavior, underscoring the potential of LHP methods in better aligning LLMs with educational objectives and fostering effective learning experiences.
 1 Introduction
-Large Language Models (LLMs) are an integral part of the modern digital landscape, and ensuring their alignment to be honest, harmless, and helpful - the ‘3H’ principle - is critical to their effective usage [1].
-However, the needs of a student learning a subject or preparing for an exam differ substantially from those of a casual internet user.
-While the latter might seek immediate answers to their queries, the former benefits more from a structured, guided approach that encourages active learning [2] and critical thinking [3].
-Therefore, if we aim to leverage LLMs to assist students’ learning, we must redefine the ‘helpfulness’ criterion in a way that specifically supports this demographic [4].
+As Large Language Models (LLMs) become integral to the digital landscape, aligning them with the ‘3H’ principle - honesty, harmlessness, and helpfulness - is crucial [1].
+In an educational context, ‘helpfulness’ takes on a new meaning; rather than providing immediate answers, LLMs should guide students through a structured step-by-step problem-solving processes to encourage active learning and critical thinking [2, 3, 4].
+This objective, which we term ‘pedagogical alignment,’ involves decomposing complex problems and providing constructive feedback and hints [5, 6].
+Previous adaptation of LLMs for education, such as the CLASS framework [7], relied on supervised fine-tuning (SFT).
+However, SFT doesn’t explicitly favor pedagogically effective responses, a key aspect of pedagogical alignment [8, 9].
 Figure 1: The image depicts the comparison between a traditional Large Language Model (LLM) interaction (left) and a pedagogically-aligned LLM interaction (right).
@@
 This elucidates the concept of pedagogical alignment, emphasizing the transformation from direct problem-solving to a guided, scaffolded learning experience.
-In this paper, we propose a novel concept of pedagogical alignment for LLMs which aligns LLMs to function as scaffolding tools [5], breaking complex problems into manageable subproblems and guiding students towards the final answer through constructive feedback and hint [6].
-This approach redefines the ‘helpfulness’ criterion of the ‘3H’ principle, tailoring it to the unique needs and learning objectives of students.
-We argue that to truly assist students, LLMs should not merely provide immediate answers, but rather guide learners through a structured, step-by-step process that encourages active learning and critical thinking.
-The pedagogical alignment of LLMs is not just a theoretical construct; it is grounded in the belief that learning can be significantly improved when students are given the opportunity to actively participate in their learning process [7] guided by feedback [8].
-Our work builds upon the framework of Conversational Learning with Analytical Step-by-Step Strategies (CLASS) [9, 10], a tool that applies above principles of learning sciences to construct an Intelligent Tutoring System (ITS) using LLMs.
-However, their approach is implemented through supervised fine-tuning of the LLM, which we identify as a limitation.
-We argue that the desired behavior to foster learning can be more effectively modeled using reinforcement learning.
-However, the challenge lies in the absence of a reward dataset necessary for reinforcement learning in education context.
-To address this challenge, we reinterpret the CLASS framework from a pedagogical alignment perspective, proposing that it can be viewed as a potential reinforcement learning problem where the rewards are associated with pedagogical alignment.
-This perspective allows us to construct a preference (reward) dataset, laying the groundwork to enhance the alignment of LLMs to better serve educational purposes.
-Next, using this reward dataset, we study the effective of three RLHF algorithms: Direct Preference Optimization (DPO) [11], Identity Preference Optimization (IPO) [12], and Kahneman Tversky Optimization (KTO) [13] for pedagogical alignment of two state-of-the-art open-source LLMs - Mistral [14] and Zephyr [15].
-We measure the the efficacy of these algorithms on pedagogical alignment by an accuracy metric.
-This metric quantifies the alignment by considering whether the correct pedagogical action was taken by the model - a crucial aspect of our pedagogical alignment approach.
-Our findings reveal that DPO and KTO significantly enhance the performance of both Mistral and Zephyr models, improving our pedagogical alignment accuracy by 37 and 23 points respectively, compared to CLASS’ SFT approach.
-Also, we examine the influence of the beta hyperparameter on the performance of these alignment algorithms since beta is a critical parameter that significantly affects the outcome of the alignment process.
-Our contributions in this paper are threefold:
-  1. 1.
-    We introduce the concept of “pedagogical alignment,” a new approach to align LLMs in a manner that fosters optimal student learning.
-  2. 2.
-    We reinterpret the CLASS framework from a pedagogical alignment perspective, which allows us to create a reward dataset.
-  3. 3.
-    Using the reward dataset, we demonstrate the effectiveness of reinforcement learning-based alignment algorithms on state-of-the-art LLMs.
-Through this work, we aim to contribute to the effective use of LLMs in education, highlighting their potential as productive learning aids and enhancing the learning experience for students.
-Table 1: Inference prompt of CLASS framework.
-Prompt: Instructions to act as a Tutorbot, an AI-powered chatbot designed to help student with a question.
-Your goal is to break problem into  subproblems for the student, and walk the student to the final answer using the subproblems and hints.
-For each student response, think which category your response falls in, and then frame the reply using this json structure:
-{ "Evaluation": "a,b,c,d,e,f,g",
-   "Action Based on Evaluation": "1,2,3,4,5,6,..,12",
-   "Subproblem State": "x,y,z",
-   "Subproblem": "..",
-   "Tutorbot": "..", }
+In this paper, we address this limitation by modeling pedagogical alignment as a learning from human preferences (LHP) problem, where desired alignment is represented by preferences.
+A major challenge in this approach is the absence of a preference dataset tailored for the educational context.
+To overcome this, we devise a method for generating a synthetic preference dataset, extending the conversational dataset generation pipeline used in CLASS [7].
+Next, using this preference dataset, we study the effectiveness of three LHP algorithms: Direct Preference Optimization (DPO) [10], Identity Preference Optimization (IPO) [11], and Kahneman Tversky Optimization (KTO) [12] for pedagogical alignment of two state-of-the-art open-source LLMs - Mistral [13] and Zephyr [14].
+We evaluate the LLMs using pedagogical alignment accuracy and F1 score which measure if the models produce the desired pedagogical responses.
+Our results show that DPO and KTO significantly improve Mistral and Zephyr’s performance, improving pedagogical alignment accuracy by 37.0 and 23.0 points, respectively, compared to SFT.
+To further investigate the effectiveness of LHP over SFT, we introduce a second preference dataset generation technique.
+This technique allows us to conduct a perplexity analysis, examining the LLM’s predisposition to provide hints and guidance versus giving direct answers.
+This analysis is critical in understanding how well the LLM aligns with our definition of pedagogical alignment, which emphasizes providing step-by-step guidance over direct answers.
+We used this technique to generate pedagogically misaligned actions, which we then compared to the aligned actions in terms of perplexity.
+Our analysis revealed that base models are more inclined to provide direct solutions than offer hints and guidance.
+While SFT partially corrects this behavior, it doesn’t fully optimize it.
+However, LHP dramatically optimizes this behavior, promoting guidance over direct solutions.
+This shift in behavior underscores the effectiveness of LHP in aligning LLMs with educational objectives, providing compelling evidence for its superiority over SFT in fostering effective learning experiences.
+2 Related Work
+Prompt: Your goal is to create a mock conversation between Student and a Tutorbot, an AI-powered chatbot designed to help Students with a question:
+"Student": "Q. {problem}",
+"Evaluation of Student Response": "..",
+"Action Based on Evaluation": "..",
+"Subproblem State": "..",
+"Tutorbot": "Lets break the problem into subproblems and tackle the subproblems one by one. Lets begin with the first subproblem...",
 Evaluation of Student Response:
-a) Correct Response
-b) Incorrect Response
-c) Partially Correct Response
-d) Ambiguous or Unclear Response
-e) ...
-Tutorbot Actions:
+a) Evaluating Incorrect Responses
+b) Evaluating Correct Responses
+c) Evaluating Partially Correct Responses
+d) Evaluating Ambiguous or Unclear or Short Responses
+e) Redirecting Off-topic Responses
+f) Responding to Student Inquiries
+g) N/A
+Prompt continues:
 If "a" is the evaluation, then:
-1) Confirm the correct answer. Check for completeness of the answer. If solution is incomplete, notify the student to complete the solution.
-Prompt continues: If "b" is the evaluation, then: 2) Notify the student about the mistake, Provide constructive feedback to pinpoint the errors, Offer hints 3) Step in to provide a solution if the student is unable to answer even after multiple attempts.
-If "c" is the evaluation, then: 4) Acknowledge the accurate parts, Promptly notify the student about the mistake, Provide constructive feedback to pinpoint the errors, Offer helpful hints 5) Step in to provide a solution if the student is unable to answer after multiple attempts If "d" is the evaluation, then: 6) Actively seek clarification through relevant follow-up questions.
-Request student to provide specific information Subproblem State: x) One of the subproblems is currently being solved y) Subproblem finished, move to next subproblem z) Subproblem finished, no next subproblem, problem finished
-2 CLASS through the lens of pedagogical alignment
-This section presents the concept of pedagogical alignment and uses it to examine the CLASS framework.
-This examination highlights the potential of application of reinforcement learning and paves the way for constructing a reward dataset to enhance the pedagogical alignment of LLMs.
-2.1 Pedagogical alignment
-In the landscape of LLMs, the term “alignment" is a concept used to describe the process of adapting LLMs to align with human preferences.
-In the context of online conversational agents, alignment specifically refers to making an LLM honest, harmless, and helpful to a typical internet user [1].
-However, the needs of a student are distinct from those of a regular internet user.
-As illustrated in Figure 1, when an internet user poses a question to a LLM, the LLM’s response should directly address the user’s query.
-This is the primary use case for out-of-the-box LLMs like Mistral [14] and Llama [16], which are designed for such interactions.
-However, to promote optimal student learning, LLMs should guide the student to reach the correct answer through their own reasoning by employing effective scaffolding [17], involving problem decomposition, constructive feedback, and timely hints.
-Here, we introduce our unique perspective of viewing the CLASS framework in the context of pedagogical alignment.
-The CLASS tutorbot differentiates itself from conventional out-of-the-box LLMs through its approach to student queries.
-Instead of directly providing answers, the CLASS tutorbot takes on a more interactive and guided approach.
-It begins by breaking the problem into smaller, manageable subproblems.
-Each subproblem presents a learning opportunity for the student, guiding them towards the final solution in a step-by-step manner.
-The CLASS framework achieves effective scaffolding through a defined discrete action and state space.
-For every student response, the CLASS tutorbot determines whether the student’s response is correct, incorrect, or ambiguous.
-For instance, if the response is incorrect, an out-of-the-box LLM’s instinctual response would be to provide the correct answer to the user.
-However, the CLASS tutorbot’s initial reaction is to offer constructive feedback to the student and allow the student to arrive at the right answer independently.
-If the student is unable to do so, then the CLASS’s eventual response is to step in and provide the correct answer (refer to action 2 and 3 in inference prompt  1).
-Similarly, if the student’s response is partially correct, the CLASS tutorbot’s first reaction is to inform the student about the correct aspects and then highlight the errors, encouraging the student to correct their answer independently (action 4 in prompt 1), as opposed to providing the complete answer, which is a typical reaction of an out-of-the-box LLM.
-Pedagogical alignment: We define the term ‘pedagogical alignment’ to describe this shift in behavior that CLASS introduces, in contrast to an out-of-the-box LLM, to promote learning [18].
-While out-of-the-box LLMs are designed to be helpful to general users, pedagogically aligned LLMs are tailored to foster optimal learning experiences for students.
-2.2 Reinforcement learning algorithms for alignment
-While the CLASS framework used SFT for pedagogical alignment, recent developments in the field of Reinforcement Learning from Human Feedback (RLHF) present promising alternatives.
-RLHF refines LLM’s behavior based on human-derived preferences or feedback, promoting a more profound congruence with human ethical standards and objectives.
+Action 1) Promptly notify the student about the mistake, Provide constructive feedback to pinpoint the errors, Offer helpful hints
+Action 2) Step in to provide a solution if the student is unable to answer even after multiple attempts.
+If "c" is the evaluation, then:
+Action 4) Acknowledge the accurate parts, Promptly notify the student about the mistake, Provide constructive feedback to pinpoint the errors, Offer helpful hints
+Action 5) Step in to provide a solution if the student is unable to answer after multiple attempts
+Subproblem states=:
+x) One of the subproblems is currently being solved
+y) Subproblem finished, moving to next subproblem
+Table 1: Synthetic conversational data generation prompt of CLASS framework.
+While CLASS introduced actions for interpretability, in our work we utilize these actions in a novel way to construct a preference dataset that distinguishes pedagogically preferred responses from less effective ones for training our LHP models.
+2.1 Algorithms for LHP
+Recent developments in the field of learning from human preferences (LHP) and reinforcement learning through human feedback (RLHF) present promising alternatives for pedagogical alignment.
+LHP/RLHF refines LLM’s behavior based on human-derived preferences or feedback, promoting a more profound congruence with human ethical standards and objectives.
 The seminal work by [9] introduced the concept of leveraging human preferences to steer the fine-tuning process of language models, thereby laying the foundational stone for RLHF.
 Subsequent advancements [15, 16] have led to notable enhancements in the RLHF pipeline, augmenting alignment efficiency and overall model performance.
-The RLHF methodology unfolds in a structured three-step process:
-  1. 1.
-    Supervised Fine-Tuning (SFT): Initially, the LLM undergoes SFT using a task-specific dataset.
-This step adjusts the model’s initial parameters to align with the downstream task’s objectives, providing a tailored foundation for further refinement.
-  2. 2.
-    Preference Modeling: Following SFT, the model generates multiple responses under varied generation parameters.
-Human evaluators then rank these responses based on their preferences.
-This ranking process informs the development of a preference model, which learns to assign scores to LLM outputs, effectively translating human judgments into quantifiable metrics.
-  3. 3.
-    Online Reinforcement Learning: Utilizing the preference model, LLMs are further aligned through online reinforcement learning algorithms such as Proximal Policy Optimization (PPO) [22].
-This phase iteratively refines the model’s parameters to maximize the alignment with human feedback, enhancing the model’s performance and ethical alignment.
 More recently, Direct Preference Optimization (DPO) [10] has emerged as a streamlined and robust advancement over RLHF, offering superior performance by forgoing the necessity for a reward model.
@@
 Then, DPO’s optimization problem can be written as:
-$L_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)\right],$ (1)
+$L_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)\right]$ (1)
 where $\sigma$ is the sigmoid function, and $\beta$ is an algorithmic parameter.
@@
 All the three methodologies, DPO, IPO, and KTO, require a dataset composed of preferences, denoted as ${(x,y_{w},y_{l})}$.
-The assembly of such datasets in a real-world context is notably labor-intensive and financially burdensome.
-However, in the subsequent section, we introduce our innovative reward/preference data generation approach, which effectively circumvents these challenges, paving the way for efficient and cost-effective pedagogical alignment of LLMs.
+The curation of such datasets in a real-world context is notably labor-intensive and financially burdensome.
+However, in the subsequent section, we introduce our innovative reward/preference data generation approach, which effectively circumvents these challenges, paving the way for cost-effective pedagogical alignment of LLMs.
+2.2 Synthetic Student Data Generation
+To create our preference dataset, we adapt the CLASS framework [7] that uses GPT-4 [17] to generate synthetic conversational datasets.
+These datasets simulate a range of interactions between students and GPT-4 using various prompts.
+An example prompt is illustrated in Table 1.
+The prompt instructs the model to first evaluate the student response (Evaluation of Student Response), consider which action to employ (Action based on Evaluation), determine the sub-problem (Subproblem State), and conditioned on these fields generate the reply to the student.
+While the CLASS framework introduced these fields to enhance interpretability for teachers and other educational stakeholders, we utilize them differently to construct a preference dataset, as we will discuss in the next section.
 3 Preference Data Generation for Pedagogical Alignment
-Figure 2: This illustration shows the process of generating pedagogically-aligned preference data using the CLASS framework.
+Figure 2: This figure shows the process of generating pedagogically-aligned preference data using the CLASS framework.
 The GPT-student asks a question, to which both the GPT-tutor and SFT-tutor respond.
@@
 In this example, the GPT-tutor’s response is deemed more pedagogically aligned because it encourages the student to engage in critical thinking and attempt the problem again, instead of directly providing the correct answer.
-This action mismatch between the tutor responses allows us to construct a preference dataset that distinguishes between the pedagogically preferred (chosen) and less effective (rejected) responses.
-The preference dataset serves as the foundation for applying reinforcement learning-based alignment algorithms, with the goal of enhancing the pedagogical alignment of large language models.
-In this section, we describe our DPO, IPO, and KTO compatible data generation pipeline.
+This action mismatch between the two tutor responses allows us to construct a preference dataset that distinguishes between the pedagogically preferred (chosen) and less effective (rejected) responses.
+In this section, we describe our DPO, IPO, and KTO compatible data generation pipeline which we also outline in figure 2.
 The objective is to generate a preference dataset $\mathcal{D}_{p}$ which necessitates distinguishing between pedagogically chosen and rejected tutor responses based on their alignment with desired scaffolding strategies.
@@
 The chosen contains the pedagogically preferred or optimal tutor responses to the context, while rejected includes those tutor responses considered less effective or misaligned with the scaffolding approach.
-CLASS Conversational Dataset $\mathcal{D}_{c}$ and Agents: Our base dataset is the CLASS conversational dataset represented by $\mathcal{D}_{c}=\{(\mathbf{x}_{i},\mathbf{y}_{i})|i=1,\ldots,N\}$, comprising $N$ student-tutor conversational turns.
+Conversational Dataset $\mathcal{D}_{c}$ and Agents: We use the conversational prompt listed in Table 1 to generate a dataset represented by $\mathcal{D}_{c}=\{(\mathbf{x}_{i},\mathbf{y}_{i})|i=1,\ldots,N\}$, comprising $N$ student-tutor conversational turns.
 Each $\mathbf{x}$ represents a sequence of student utterances, and each corresponding $\mathbf{y}$ represents the tutor response.
@@
 The conversation starts with GPT-student $\mathcal{A}_{\text{G}}$ posing a question ($Q$) sourced from a question bank to which both GPT-tutor $\mathcal{A}_{\text{T}}$ and SFT-tutor $\mathcal{A}_{\text{S}}$ provide responses in the following JSON structure:
-$\mathbf{y}_{i}=\left\{\begin{array}[]{l}\text{``Evaluation of Student Response": ``a, b, c, \ldots",}\\ \text{``Action Based on Evaluation": ``1, 2, 3, \ldots, 12",}\\ \text{``Subproblem State": ``x, y, z",}\\ \text{``Subproblem": ``...",}\\ \text{``Tutorbot": ``..." }\end{array}\right.$ (2)
+$\mathbf{y}_{i}=\left\{\begin{array}[]{l}\text{``Eval of Student Response": ``a, b, c, \ldots",}\\ \text{``Action Based on Eval": ``1, 2, 3, \ldots, 12",}\\ \text{``Subproblem State": ``x, y, z",}\\ \text{``Subproblem": ``...",}\\ \text{``Tutorbot": ``..." }\end{array}\right.$ (2)
 Defining Context: We define the context $C_{t}$ for any given response at time $t$ as all preceding conversational turns up to that point:
@@
 […]
-$\text{and }f(\mathbf{y}_{t}^{\mathcal{A}_{T}})\neq f(\mathbf{y}_{t}^{\mathcal{A}_{S}}),\forall t\in\{1,\ldots,N\}\big{\}}$ (4)
 where $C_{t}=x,y_{w}=y_{choosen},y_{l}=y_{rejected}$ from DPO equation (1) from section 2.1, and the function $f$ is the key to distinguishing between chosen and rejected tutor responses.
 $f$ indicates a discrepancy between $\mathcal{A}_{\text{T}}$’s and $\mathcal{A}_{\text{S}}$’s responses if any of the fields “Evaluation", “Action Based on Evaluation", or “Subproblem State" diverge between $\mathcal{A}_{\text{T}}$ and $\mathcal{A}_{\text{S}}$, $f$ asserts $\mathcal{A}_{\text{T}}$’s response as preferable, under the assumption that $\mathcal{A}_{\text{T}}$ offers a more accurate pedagogical model, alongside the rejected response from $\mathcal{A}_{\text{S}}$.
+Figure 2 illustrates this conception with a concrete example: GPT-tutor ($\mathcal{A}_{\text{T}}$) and SFT-tutor ($\mathcal{A}_{\text{S}}$) provide different responses since the Action fields are different, so we choose the GPT-tutor response as the accepted one and the SFT-tutor response as the rejected one.
 This condition is captured through the function $f$—mapping responses to their pedagogical attributes:
-$f(\mathbf{y}_{i}^{\mathcal{A}})=\left(\mathrm{Eval}(\mathbf{y}_{i}^{\mathcal{A}}),\,\mathrm{ActionEval}(\mathbf{y}_{i}^{\mathcal{A}}),\,\mathrm{SubprobState}(\mathbf{y}_{i}^{\mathcal{A}})\right)$ (5)
-where,
+${f(\mathbf{y}_{i}^{\mathcal{A}})=\left(\mathrm{Eval}(\mathbf{y}_{i}^{\mathcal{A}}),\mathrm{Action}(\mathbf{y}_{i}^{\mathcal{A}}),\mathrm{Subprob}(\mathbf{y}_{i}^{\mathcal{A}})\right)}$ (5)
   • $\text{Eval}(\mathbf{y}_{i}^{\mathcal{A}})$ extracts insights on the tutor’s grading of student’s solution or question understanding.
-  • $\text{ActionEval}(\mathbf{y}_{i}^{\mathcal{A}})$ outlines the tutor’s recommended actions based on the evaluation.
-  • $\text{SubprobState}(\mathbf{y}_{i}^{\mathcal{A}})$ provides an understanding of the tutor’s perception of the student’s current understanding or progress on sub-problems within a larger problem context.
+  • $\text{Action}(\mathbf{y}_{i}^{\mathcal{A}})$ outlines the tutor’s recommended actions based on the evaluation.
+  • $\text{Subprob}(\mathbf{y}_{i}^{\mathcal{A}})$ provides an understanding of the tutor’s perception of the student’s current understanding or progress on sub-problems within a larger problem context.
 The evaluation of these attributes by $f$ facilitates an objective basis to deem one response as chosen (preferable) and another as rejected (less aligned with desired pedagogical outcomes) through a systematic assessment of their pedagogical value and alignment.
 4 Experiments
-In this section, we provide a comprehensive overview of our dataset construction, experimental design, and the subsequent findings derived from the application of state-of-the-art RLHF algorithms to train pedagogically-aligned LLMs.
+In this section, we provide a comprehensive overview of our dataset construction, experimental design, and the subsequent findings derived from the application of state-of-the-art LHP algorithms to train pedagogically-aligned LLMs.
 4.1 Dataset and Evaluation
-Our experimental design requires an extensive conversational dataset to train and test SFT and RLHF models effectively.
-The CLASS framework [9] used GPT-4 to generate 610 conversations to train their SFT model.
-However, we need to generate a larger volume of conversations to train RLHF models and conduct comprehensive testing.
+Our experimental design requires an extensive conversational dataset to train and test SFT and LHP models effectively.
+The CLASS framework used GPT-4 to generate 610 conversations to train their SFT model.
+However, we need to generate a larger volume of conversations to train LHP models and conduct comprehensive testing.
 Following the strategy outlined by CLASS, we set about creating these additional conversations.
-We used the CLASS scaffolding prompts to generate the problems, subproblems, and associated hints, which were based on biology topics from the OpenStax college book [23], consistent with CLASS’ context.
+We used the CLASS scaffolding prompts to generate the problems, subproblems, and associated hints, which were based on biology topics from the OpenStax college books [18], consistent with CLASS’ context.
 These problems seed the context of the conversations generated through GPT-4 in the next step using the CLASS conversational prompt.
@@
 Each conversation, on average, comprises approximately 8.2 rounds and 1311.0 words.
-This comprehensive dataset provides a solid platform for training and comparing SFT and RLHF models.
-To create the training and testing set for SFT and RLHF, we randomly sampled our conversational dataset into 3 partitions: the SFT training dataset $D_{c}$ (600 conversations), the RLHF training dataset $D_{p}$ (600 conversations), and the test dataset $D_{t}$ (450 conversations).
+This comprehensive dataset provides a solid platform for training and comparing SFT and LHP models.
+To create the training and testing set for SFT and LHP, we randomly sampled our conversational dataset into 3 partitions: the SFT training dataset $D_{c}$ (600 conversations), the LHP training dataset $D_{p}$ (600 conversations), and the test dataset $D_{t}$ (450 conversations).
 As each round within the conversations is processed into a separate data point, $D_{c}$ contains 4942 data points, $D_{p}$ contains 4921 data points, and $D_{t}$ contains 3701 data points.
 For the evaluation, we focus on the three classification fields within the model JSON responses described in Section 3: Evaluation of Student Response, Action Based on Evaluation, and Subproblem State.
-We employ accuracy in these three fields to measure the pedagogical alignment of LLMs since these metrics offer a comprehensive view of the model’s performance, evaluating the correctness of the responses, the appropriateness of the actions, and the state of problem-solving.
-Hence, LLMs that achieve higher accuracy in these fields are considered to provide more pedagogically aligned student assistance.
+We employ accuracy and F1 score in these three fields to measure the pedagogical alignment of LLMs since these metrics offer a comprehensive view of the model’s performance in evaluating the correctness of the responses (Evaluation of Student Response), the appropriateness of the actions (Action Based on Evaluation), and the state of problem-solving (Subproblem State).
+Hence, LLMs that achieve higher accuracy and F1 score in these fields are considered to provide more pedagogically aligned assitance to students.
+Model Metric SFT DPO IPO KTO
+Mistral Acc 35.0 (34,34,37) 74.0 (72,72,78) 57.7 (56,56,61) 75.3 (73,73,80)
+F1 0.34 (0.42,0.24,0.37) 0.56 (0.63,0.37,0.67) 0.48 (0.56,0.33,0.56) 0.55 (0.62,0.37,0.65)
+Zephyr Acc 10.7 (12,11,9) 74.0 (72,71,79) 21.7 (22,22,21) 74.7 (73,72,79)
+F1 0.14 (0.21,0.12,0.10) 0.56 (0.65,0.38,0.66) 0.26 (0.32,0.19,0.28) 0.54 (0.63,0.37,0.63)
 Table 2: SFT vs DPO, IPO, KTO.
-Each cell shows the average accuracy across three classification fields - Evaluation of student response, Action based on the evaluation, and Subproblem State, respectively.
-We observed that for both Mistral (Mistral-7B-Instruct-v0.2) and Zephyr (zephyr-7b-beta), all three RL alignment algorithms - DPO, KTO, and IPO - showed notable accuracy gains in accuracy across all fields as compared to SFT.
-For both models, DPO and KTO outperformed IPO.
-IPO performed best for Vicuna (vicuna-7b-v1.5) compared to the other two.
-Model SFT DPO IPO KTO
-Mistral 35.0 (34,34,37) 74.0 (72,72,78) 57.7 (56,56,61) 74.7 (73,73,80)
-Vicuna 74.3 (73,70,80) 75.0 (73,72,80) 76.0 (74,72,82) 75.3 (73,72,81)
-Zephyr 10.7 (12,11,9) 74.0 (72,71,79) 21.7 (22,22,21) 74.7 (73,72,79)
-Table 3: LLM Performance on key benchmarks.
-This table presents the performance of Mistral (Mistral-7B-Instruct-v0.2), Zephyr (zephyr-7b-beta), and Vicuna(vicuna-7b-v1.5) on different NLP benchmarks [24, 25].
-We selected Mistral and Zephyr for our experiments based on their superior performance across various benchmarks, which provided a thorough and fair assessment across a spectrum of reasoning and general knowledge tasks.
-We chose Vicuna because it served as the base model for the CLASS framework.
-Model Average ARC HellaSwag MMLU TruthfulQA Winogrande GSM8K
-Mistral 65.71 63.14 84.88 60.78 68.26 77.19 40.03
-Zephyr 61.95 62.03 84.36 61.07 57.45 77.74 29.04
-Vicuna 52.06 53.24 77.39 50.82 50.33 72.06 8.11
+Each cell shows the average accuracy and average F1 score across three classification fields - Evaluation of student response, Action based on the evaluation, and Subproblem State, respectively.
+We observed that for both Mistral (Mistral-7B-Instruct-v0.2) and Zephyr (zephyr-7b-beta), all three alignment algorithms - DPO, KTO, and IPO - showed notable gains in accuracy and F1 score across all fields as compared to SFT.
+Across both models, DPO and KTO exhibited comparable performance, outperforming IPO.
 4.2 Models and and Training Details
-We experimented on 3 different LLMs: Mistral (Mistral-7B-Instruct-v0.2), Zephyr (zephyr-7b-beta), and Vicuna (vicuna-7b-v1.5).
-We selected Mistral and Zephyr based on the their top performance on diverse NLP benchmarks (Table 3).
-Among all, several benchmark in particular evaluate scientific knowledge and education-based question answering capabilities: ARC (multiple choice questions from grade school science exams) [26], MMLU (multiple choice questions from 57 high school and college subjects across STEM, the humanities, the social sciences, and more) [27], and GSM8K (linguistically diverse grade school math word problems) [28].
-These benchmarks offer a comprehensive, unbiased evaluation on the models’ performance, especially in education setting.
-Furthermore, despite Vicuna’s inferior performance compared to Mistral and Zephyr, we included it in our selection because it was the base model for the CLASS framework’s initial experiments and represented the state of the art at its launch.
-Our experiments included training and evaluating SFT and RLHF models.
-It took around 22 NVIDIA A6000 GPU hours for one cycle (SFT + Evaluation + RLHF + Evaluation).
+We experimented on two different LLMs: Mistral (Mistral-7B-Instruct-v0.2) and Zephyr (zephyr-7b-beta).
+We selected Mistral and Zephyr based on their top performance on diverse NLP benchmarks on OpenLLM leaderboard [19, 20] which evaluate scientific knowledge and education-based question answering capabilities: ARC (reasoning based science questions) [21], MMLU (multiple choice questions from 57 high school and college subjects across STEM, the humanities, the social sciences, and more) [22], and GSM8K (grade school math word problems) [23].
+These benchmarks offer a comprehensive evaluation on the models’ performance, especially in education setting.
+Our experiments included training and evaluating SFT and LHP models.
+It took around 22 NVIDIA A6000 GPU hours for one cycle (SFT + Evaluation + LHP + Evaluation).
 A total of 17 cycles, which are 374 GPU hours, are needed to complete all the experiments.
@@
 For the choice of SFT hyperparameters, we refered to parameters used to instruct fine-tune Vicuna [24]: We employed a learning rate of 2e-5 with the AdamW optimizer [25], a batch size of 16, a cosine scheduler, a weight decay of 0.05, a warmup ratio of 0.1, and a total number of 3 epochs.
-We conducted the RLHF experiments using TRL [31].
-For the choice of RLHF hyperparamters, we referred to parameters used to align Zephyr [32]: We employed a learning rate of 1e-7 with the AdamW optimizer, a batch size of 16, a cosine scheduler, a weight decay of 0.05, a warmup ratio of 0.1, and a total number of 3 epochs.
-4.3 Experiment 1: SFT vs DPO, IPO, and KTO
+We conducted the LHP experiments using TRL [26].
+For the choice of LHP hyperparamters, we referred to parameters used to align Zephyr [27]: We employed a learning rate of 1e-7 with the AdamW optimizer, a batch size of 16, a cosine scheduler, a weight decay of 0.05, a warmup ratio of 0.1, and a total number of 3 epochs.
+4.3 Main findings: SFT vs LHP
 In this experiment, we compared three alignment algorithms: Direct Preference Optimization (DPO), Identity Preference Optimization (IPO), and Kahneman Tversky Optimization (KTO) on the SFT models.
-We measured the accuracy across three classification fields in the responses: Evaluation of Student Response, Action based on Evaluation, and Subproblem State.
-We used the same RLHF training setting as described in Section 4.2 with a beta of 0.1 for DPO, IPO, and KTO across all models.
-After applying the alignment algorithms, we observed significant accuracy improvements on Mistral and Zephyr SFT models.
-Specifically, Mistral demonstrated improvements of 39%, 22.7%, and 39.7% from DPO, IPO, and KTO respectively, while Zephyr showed improvements of 63.3%, 11%, and 64% respectively (Table 2).
+We measured the accuracy and F1 score across three classification fields in the responses: Evaluation of Student Response, Action based on Evaluation, and Subproblem State.
+We used the same LHP training setting as described in Section 4.2 with a beta of 0.1 for DPO, IPO, and KTO across all models.
+After applying the alignment algorithms, we observed significant performance improvements on Mistral and Zephyr SFT models.
+Specifically, Mistral demonstrated accuracy increase of 39%, 22.7%, and 39.7% from DPO, IPO, and KTO respectively, while Zephyr showed improvements of 63.3%, 11%, and 64% respectively (Table 2).
+We observed similar increase in F1 score as well: Mistral showed increase of 0.22, 0.14, and 0.21 from DPO, IPO, and KTO, and Zephyr showed improvements of 0.42, 0.12, and 0.4 respectively.
 Notably, for both Mistral and Zephyr, DPO and KTO significantly outperformed IPO.
-However, the Vicuna model, already noted for its high SFT performance, showed only marginal improvements from the alignment; all three algorithms exhibited roughly similar performance.
 Our results are in alignment with the findings of similar studies [27], which have reported superior performance of DPO and KTO over IPO on Zephyr when the model was fine-tuned on preference datasets, such as Intel’s ORCA dataset pairs [28] and the Ultra-Feedback dataset [14].
-This experiment underscored the pivotal role of alignment algorithms in improving the accuracy of SFT models.
-Mistral and Zephyr, the top performing models, exhibited pronounced improvements over Vicuna.
-Therefore, we demonstrated the efficacy of pedagogical alignment on state of the art models.
-4.4 Experiment 2: Effect of Beta on DPO and IPO
+This experiment underscored the pivotal role of alignment algorithms in improving the accuracy and F1 score of SFT models.
+Mistral and Zephyr, the top performing models, exhibited pronounced improvements using LHP algorithms compared to SFT.
+4.4 Examining LHP through lens of perplexity
+For understanding the effectiveness of LHP algorithms compared to SFT, we propose a second preference dataset generation technique.
+This technique specifically allows us to perform a perplexity analysis to compare the LLM’s tendency to provide hints and guidance versus direct responses.
+The essence of pedagogical alignment is the LLM’s propensity to offer guidance (as seen in Actions 1 and 4) over direct solutions (as in Actions 2 and 5).
+By this definition, Action 2 can never precede Action 1, and likewise, Action 5 can never precede Action 4 (for more details about actions please refer to Table 1).
+With this critical insight, we use GPT-4 to generate pedagogically misaligned Actions (Action 2 for the first occurrence of Action 1 and Action 5 for the first occurrence of Action 4).
+This approach allows us to compare the perplexity of responses corresponding to Action 1 and its misaligned counterpart (Action 2), and similarly for Actions 4 and 5.
+Building upon the test split (450 conversations) of the originally generated dataset, we generate an alternative opposite action (Action 2 for Action 1, and Action 5 for Action 4) for the first occurrence of Action 1 or 4 in each conversation.
+We then use this newly constructed Action dataset to measure the perplexity of responses from the original instruct model (Base), its SFT fine-tuned counterpart (SFT), and its LHP version (DPO) (Table 3).
+Our analysis reveals that for both Mistral and Zephyr Base models, responses with Action 1 have higher perplexity than those with Action 2, and likewise for Actions 4 and 5.
+This suggests that the base models are more inclined to provide direct solutions than offer hints and guidance.
+The SFT process adjusts this behavior by increasing the perplexity of responses with Actions 2 and 5, thus making the models more likely to produce responses with Actions 1 and 4.
+However, SFT alone doesn’t fully optimize this behavior.
+As seen in the perplexity results for Mistral and Zephyr SFT models (Table 3), there is only a minimal decrease or even a slight increase in the perplexity of Action 1 and 4.
+This indicates that SFT, while useful, doesn’t significantly shift the model’s behavior towards providing more guidance over direct solutions.
+In contrast, LHP dramatically optimizes this behavior.
+For the Mistral model, the perplexity of Action 1 and 4 decreases from 2.05 and 2.14 to 1.89 and 1.96 (SFT vs DPO) respectively.
+Similarly, for the Zephyr model, the perplexity of Action 1 and 4 decreases from 2.09 and 2.18 (SFT) to 1.97 and 2.05 (DPO) respectively.
+This decrease in perplexity indicates a higher likelihood of the model providing guidance over direct solutions.
+Simultaneously, the perplexity of Action 2 and 5 increases and from 13.9 and 13.06 (Zephyr SFT) to 14.9 and 14.38 (Zephyr DPO).
+This increase in perplexity indicates a lower likelihood of the model providing direct solutions, further emphasizing the model’s shift towards providing guidance.
+Model A1 A2 A4 A5
+Mistral (Base) 2.11 2.06 2.00 1.97
+Mistral (SFT) 2.05 5.55 2.14 5.40
+Mistral (DPO) 1.89 5.48 1.96 5.34
+Zephyr (Base) 2.05 2.09 2.03 2.07
+Zephyr (SFT) 2.09 13.9 2.18 13.06
+Zephyr (DPO) 1.97 14.9 2.05 14.38
+Table 3: Perplexity of responses.
+Perplexity analysis reveals that the base models of Mistral and Zephyr are more inclined to provide direct solutions (Action 2 and 5) than guidance and hints (Action 1 and 4).
+The SFT process partially shifts this behavior, increasing the perplexity on responses with Action 2 and 5, thus making the models more likely to produce responses with Action 1 and 4.
+However, the SFT process alone doesn’t fully optimize this behavior, as seen by the minimal decrease or slight increase in the perplexity of Action 1 and 4.
+In contrast, DPO dramatically optimizes this behavior by reducing perplexity on Action 1 and 4 while simultaneously increasing perplexity on Action 2 and 5.
+This shift in perplexity indicates a significant behavioral shift in the LLM, making it more aligned with the role of a tutorbot by favoring guidance over direct solutions.
+For more details about actions please refer to Table 1.
+4.5 Effect of Beta on DPO and IPO
 Figure 3: DPO and IPO performance as a function of beta.
-This figure depicts how the beta hyper-parameter affects the average classification accuracy of DPO and IPO aligned models.
-We found that both DPO and IPO performances were significantly influenced by beta adjustments.
-Specifically, a beta of 0.3 optimized results for DPO, with Mistral and Zephyr achieving around 75% accuracy, while a higher beta of 0.9 notably reduced their accuracy.
-In contrast, IPO performed best with a beta of 0.1, and increasing the beta decreased the accuracy.
+We found that classification accuracy and F1 score using both DPO and IPO were significantly influenced by beta hyper-parameter adjustments.
+Specifically, a beta of 0.3 optimized results for DPO, with Mistral and Zephyr achieving around 75% accuracy and 0.56 F1 score, while a higher beta of 0.9 notably reduced accuracy and F1 score.
+In contrast, IPO performed best with a beta of 0.1 and increasing beta decreased accuracy and F1 score.
 In this experiment, we analyzed the impact of the beta parameter on the performance of DPO and IPO alignment algorithms.
@@
 Nevertheless, the choice of beta depends on the specific model and dataset and thus necessitates an empirical study.
-Similar to experiments in Section 4.3, we measured the average accuracy across the three classification fields.
+Similar to experiments in Section 4.3, we measured the average accuracy and F1 score across the three classification fields.
 We found that the performance of DPO and IPO is sensitive to the choice of beta.
-As depicted in figure 3, for DPO, we observed that a beta value of 0.3 yielded the best results: Mistral and Zephyr achieved an average accuracy of 74.7% and 75%.
+As depicted in Figure 3, for DPO, we observed that a beta value of 0.3 yielded the best results: Mistral and Zephyr achieved an average accuracy of 74.7% and 75% and an average F1 score of 0.55 and 0.56.
 Conversely, a larger beta value, such as 0.9, diminished the performance of the models.
-Zephyr suffered a large loss of 25.7 % in average accuracy from using 0.9 instead of 0.3.
+Zephyr suffered a large loss of 25.7 % in average accuracy from using 0.9 instead of 0.3, and a loss of 0.13 in average F1 score.
 As opposed to IPO, we observed that a beta of 0.1 produced the best performance (Figure 3).
 A larger choice of beta in IPO almost always came with a loss of performance.
-As beta increases, Mistral’s accuracy dropped from 57.7% to 36 %, while Zephyr’s accuracy dropped from 21.7 % to 11.3 %.
+As beta increased, Mistral’s accuracy dropped from 57.7% to 36 %, and its F1 score decreased from 0.55 to 0.49, while Zephyr’s accuracy dropped from 21.7 % to 11.3 %, and its F1 score decreased from 0.56 to 0.43.
 Our results corroborate the evidence from related studies [27], demonstrating that various alignment algorithms exhibit distinct sensitivities to beta values, necessitating tailored beta selections.
 5 Conclusion
-In this paper, we have introduced and explored the concept of pedagogical alignment of LLMs to align these models with educational objectives that are aimed at fostering optimal student learning.
-By reinterpreting the CLASS framework as a reinforcement learning problem, we constructed a reward dataset, thereby providing a foundation for enhancing the pedagogical alignment of LLMs.
-Our experimental results, derived from applying reinforcement learning-based alignment algorithms on state-of-the-art open-source LLMs, corroborate our hypothesis.
-The RLHF methods we employed have proven to outperform the SFT approach significantly, favoring the alignment of LLMs with pedagogical objectives.
-Furthermore, our study highlights the potential of online feedback for enhancing the performance of pedagogically-aligned LLMs.
-In conclusion, while our research has made significant strides in aligning LLMs with educational needs, further exploration and refinement of reinforcement learning methods for this purpose remain necessary.
-We anticipate that our work will serve as a stepping stone towards the optimal integration of LLMs in educational settings, paving the way for future research in this exciting domain.
-Acknowledgments
-This work was supported by NSF grants 1842378, ONR grant N0014-20-1-2534, AFOSR grant FA9550-22-1-0060, and a Vannevar Bush Faculty Fellowship, ONR grant N00014-18-1-2047.
+In this paper, we have investigated the application of learning from human preferences (LHP) methods to align LLMs with educational objectives, aiming to foster optimal student learning.
+By constructing a preference dataset specifically designed for pedagogical alignment, we have laid the foundation for enhancing the effectiveness of LLMs in educational contexts.
+Our experimental results, derived from applying LHP alignment algorithms on state-of-the-art open-source LLMs, demonstrate the superiority of LHP methods over the traditional supervised fine-tuning (SFT) approach, significantly improving the alignment of LLMs with pedagogical objectives.
+Our another key contribution is the development of a novel approach for evaluating the pedagogical effectiveness of LLMs.
+By curating a second preference dataset that compares the perplexity of responses offering scaffolding versus those providing direct solutions, we have introduced a new methodology that quantifies the extent to which LLMs prioritize pedagogically effective actions.
+This approach opens up new possibilities for assessing and optimizing the performance of LLMs in educational settings, providing a valuable tool for researchers and practitioners alike.

### v3

@@
 Abstract
-Large Language Models (LLMs) have shown remarkable potential in various domains, including education.
-However, their application in educational contexts requires careful ‘pedagogical alignment’, where these models are tuned to assist students in a step-by-step problem-solving process, rather than providing direct answers.
-In this paper, we investigate algorithms for learning from human preferences (LHP) to achieve this alignment objective.
-A key challenge in this process is the lack of preference datasets to guide the alignment.
-To address this, we propose a novel approach for constructing such a dataset using synthetic data generation techniques.
-Leveraging this dataset, our experiments reveal the superiority of LHP methods over standard supervised fine-tuning (SFT) for pedagogical alignment.
-Applying LHP algorithms to Mistral and Zephyr, we observe a significant improvement in their ability to guide students, enhancing the pedagogical alignment accuracy by 37.0 and 23.0 points, respectively, compared to SFT.
-Additionally, we introduce a unique approach for evaluating the pedagogical effectiveness of LLMs using perplexity analysis.
-We curate a second preference dataset, comparing the perplexity of LLMs’ responses when offering scaffolded step-by-step guidance versus direct solutions.
-This allows us to quantify the extent to which LLMs prioritize pedagogically effective actions in their generation.
+Large Language Models (LLMs), when used in educational settings without pedagogical fine-tuning, often provide immediate answers rather than guiding students through the problem-solving process.
+This approach falls short of pedagogically best practices and limits their effectiveness as educational tools.
+We term the objective of training LLMs to emulate effective teaching strategies as ‘pedagogical alignment.’ In this paper, we investigate Learning from Human Preferences (LHP) algorithms to achieve this alignment objective.
+A key challenge in this process is the scarcity of high-quality preference datasets to guide the alignment.
+To address this, we propose a novel approach for constructing a large-scale dataset using synthetic data generation techniques, eliminating the need for time-consuming and costly manual annotation.
+Leveraging this dataset, our experiments with Llama and Mistral models demonstrate that LHP methods outperform standard supervised fine-tuning (SFT), improving pedagogical alignment accuracy by 13.1% and 8.7% respectively.
+Existing evaluation methods also lack quantitative metrics to adequately measure the pedagogical alignment of LLMs.
+To address this gap, we propose novel perplexity-based metrics that quantify LLMs’ tendency to provide scaffolded guidance versus direct answers, offering a robust measure of pedagogical alignment.
 Our analysis provides compelling evidence for the superiority of LHP methods over SFT in optimizing LLMs’ behavior, underscoring the potential of LHP methods in better aligning LLMs with educational objectives and fostering effective learning experiences.
+Code and models are available here.
+Equal contribution.
 1 Introduction
-As Large Language Models (LLMs) become integral to the digital landscape, aligning them with the ‘3H’ principle - honesty, harmlessness, and helpfulness - is crucial [1].
-In an educational context, ‘helpfulness’ takes on a new meaning; rather than providing immediate answers, LLMs should guide students through a structured step-by-step problem-solving processes to encourage active learning and critical thinking [2, 3, 4].
-This objective, which we term ‘pedagogical alignment,’ involves decomposing complex problems and providing constructive feedback and hints [5, 6].
-Previous adaptation of LLMs for education, such as the CLASS framework [7], relied on supervised fine-tuning (SFT).
-However, SFT doesn’t explicitly favor pedagogically effective responses, a key aspect of pedagogical alignment [8, 9].
+Large Language Models (LLMs) are increasingly being aligned with the ‘3H’ principle - honesty, harmlessness, and helpfulness (Askell et al., 2021).
+However, these definitions often require domain-specific adaptations to be truly effective.
+In educational contexts, for instance, ‘helpfulness’ takes on a new meaning; rather than providing immediate answers, LLMs should guide students through a structured step-by-step problem-solving processes to encourage active learning and critical thinking (Paul and Elder, 1995; Prince, 2004; Chi, 2009).
+This objective, which we term ‘pedagogical alignment,’ involves breaking complex problems into manageable steps and providing hints and scaffolded guidance rather than direct answers (Bransford et al., 2000; Hattie and Timperley, 2007).
+Previous adaptation of LLMs for education, such as the CLASS framework (Sonkar et al., 2023), relied on supervised fine-tuning (SFT).
+However, SFT doesn’t explicitly favor pedagogically effective responses, a key aspect of pedagogical alignment (Christiano et al., 2017; Ziegler et al., 2019).
 Figure 1: The image depicts the comparison between a traditional Large Language Model (LLM) interaction (left) and a pedagogically-aligned LLM interaction (right).
@@
 This elucidates the concept of pedagogical alignment, emphasizing the transformation from direct problem-solving to a guided, scaffolded learning experience.
-In this paper, we address this limitation by modeling pedagogical alignment as a learning from human preferences (LHP) problem, where desired alignment is represented by preferences.
-A major challenge in this approach is the absence of a preference dataset tailored for the educational context.
-To overcome this, we devise a method for generating a synthetic preference dataset, extending the conversational dataset generation pipeline used in CLASS [7].
-Next, using this preference dataset, we study the effectiveness of three LHP algorithms: Direct Preference Optimization (DPO) [10], Identity Preference Optimization (IPO) [11], and Kahneman Tversky Optimization (KTO) [12] for pedagogical alignment of two state-of-the-art open-source LLMs - Mistral [13] and Zephyr [14].
+In this paper, we propose a novel approach to achieve pedagogical alignment by modeling it as learning from human preferences (LHP).
+This method allows us to represent desired teaching behaviors as preferences, enabling more nuanced optimization than SFT.
+However, a significant challenge in applying LHP is the scarcity of high-quality, pedagogically-sound preference data.
+To address this, we introduce a key innovation: leveraging the structured output from the CLASS framework (Sonkar et al., 2023) to create meaningful preference pairs of pedagogically aligned and misaligned responses.
+This enables us to generate a large-scale, synthetic preference dataset that transforms abstract tutoring principles into concrete, comparative examples, allowing LHP algorithms to learn effective teaching strategies.
+Next, using this preference dataset, we study the effectiveness of three LHP algorithms: Direct Preference Optimization (DPO) (Rafailov et al., 2023), Identity Preference Optimization (IPO) (Azar et al., 2023), and Kahneman Tversky Optimization (KTO) (Ethayarajh et al., 2024) for pedagogical alignment of three open-source LLMs - Llama Dubey et al.
+(2024), Mistral (Jiang et al., 2023) and Zephyr (Tunstall et al., 2023).
 We evaluate the LLMs using pedagogical alignment accuracy and F1 score which measure if the models produce the desired pedagogical responses.
-Our results show that DPO and KTO significantly improve Mistral and Zephyr’s performance, improving pedagogical alignment accuracy by 37.0 and 23.0 points, respectively, compared to SFT.
-To further investigate the effectiveness of LHP over SFT, we introduce a second preference dataset generation technique.
-This technique allows us to conduct a perplexity analysis, examining the LLM’s predisposition to provide hints and guidance versus giving direct answers.
-This analysis is critical in understanding how well the LLM aligns with our definition of pedagogical alignment, which emphasizes providing step-by-step guidance over direct answers.
-We used this technique to generate pedagogically misaligned actions, which we then compared to the aligned actions in terms of perplexity.
+Our results show that DPO and KTO significantly improve Llama, Mistral and Zephyr’s performance, improving pedagogical alignment accuracy by 13.1%, 8.7% and 50.0%, respectively, compared to SFT.
+Existing evaluation methods for LLMs in education also fail to adequately quantify pedagogical alignment.
+To address this gap and compare LHP with SFT, we introduce a second preference dataset generation technique and a novel perplexity-based evaluation approach.
+This allows us to quantitatively measure an LLM’s tendency to provide step-by-step guidance versus direct answers, offering a robust metric for pedagogical alignment.
+Using our technique we generate pedagogically misaligned actions, which we then compared to the aligned actions in terms of perplexity.
 Our analysis revealed that base models are more inclined to provide direct solutions than offer hints and guidance.
@@
 Recent developments in the field of learning from human preferences (LHP) and reinforcement learning through human feedback (RLHF) present promising alternatives for pedagogical alignment.
-LHP/RLHF refines LLM’s behavior based on human-derived preferences or feedback, promoting a more profound congruence with human ethical standards and objectives.
-The seminal work by [9] introduced the concept of leveraging human preferences to steer the fine-tuning process of language models, thereby laying the foundational stone for RLHF.
-Subsequent advancements [15, 16] have led to notable enhancements in the RLHF pipeline, augmenting alignment efficiency and overall model performance.
-More recently, Direct Preference Optimization (DPO) [10] has emerged as a streamlined and robust advancement over RLHF, offering superior performance by forgoing the necessity for a reward model.
+RLHF refines LLM’s behavior based on human-derived preferences or feedback, promoting a more profound congruence with human ethical standards and objectives.
+The seminal work by (Ziegler et al., 2019) introduced the concept of leveraging human preferences to steer the fine-tuning process of language models, thereby laying the foundational stone for RLHF.
+Subsequent advancements (Nakano et al., 2021; Ouyang et al., 2022) have led to notable enhancements in the RLHF pipeline, augmenting alignment efficiency and overall model performance.
+More recently, Direct Preference Optimization (DPO) (Rafailov et al., 2023) has emerged as a streamlined and robust advancement over RLHF, offering superior performance by forgoing the necessity for an explicit reward model.
 DPO enhances alignment efficiency by directly optimizing a preference-based loss function, thereby simplifying the implementation and the operational efficiency.
@@
 Then, DPO’s optimization problem can be written as:
-$L_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)\right]$ (1)
+$W=\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)},L=\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)},\\ L_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\left[W-L\right]$ (1)
 where $\sigma$ is the sigmoid function, and $\beta$ is an algorithmic parameter.
-However, DPO has been observed to over-fit to the preference dataset, and is believed to be sensitive to hyperparameter adjustments.
-To prevent the over-fitting behaviour of DPO’s optimization algorithm, recent proposals include Identity Preference Optimisation (IPO) [11] and Kahneman-Tversky Optimization (KTO) [12].
-All the three methodologies, DPO, IPO, and KTO, require a dataset composed of preferences, denoted as ${(x,y_{w},y_{l})}$.
+DPO tends to overfit to the preference dataset and is sensitive to hyperparameter tuning.
+Recently Identity Preference Optimization (IPO) (Azar et al., 2023) was introduced to address DPO’s overfitting issues.
+With addition of a regularization term to the DPO’s loss function, IPO also trains models to convergence without requiring early stopping using:
+$L_{\text{IPO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\\ \left[\left(\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\frac{1}{2\beta}\right)^{2}\right]$
+Both DPO and IPO methodologies necessitate a dataset comprising paired preferences, denoted as ${(x,y_{w},y_{l})}$.
+The assembly of such datasets in a real-world context is notably labor-intensive and financially burdensome.
+Kahneman-Tversky Optimization (KTO) (Ethayarajh et al., 2024) presents a solution to this challenge by introducing a loss function formulated exclusively on the basis of singular instances identified as “good” or “bad”.
+All three methods, DPO, IPO, and KTO, require a dataset consisting of prompts, accepted responses, and rejected responses.
 The curation of such datasets in a real-world context is notably labor-intensive and financially burdensome.
@@
 2.2 Synthetic Student Data Generation
-To create our preference dataset, we adapt the CLASS framework [7] that uses GPT-4 [17] to generate synthetic conversational datasets.
+To create our preference dataset, we adapt the CLASS framework (Sonkar et al., 2023) that uses GPT-4 Bubeck et al.
+(2023) to generate synthetic conversational datasets.
 These datasets simulate a range of interactions between students and GPT-4 using various prompts.
@@
 The conversation starts with GPT-student $\mathcal{A}_{\text{G}}$ posing a question ($Q$) sourced from a question bank to which both GPT-tutor $\mathcal{A}_{\text{T}}$ and SFT-tutor $\mathcal{A}_{\text{S}}$ provide responses in the following JSON structure:
-$\mathbf{y}_{i}=\left\{\begin{array}[]{l}\text{``Eval of Student Response": ``a, b, c, \ldots",}\\ \text{``Action Based on Eval": ``1, 2, 3, \ldots, 12",}\\ \text{``Subproblem State": ``x, y, z",}\\ \text{``Subproblem": ``...",}\\ \text{``Tutorbot": ``..." }\end{array}\right.$ (2)
+[…]
 Defining Context: We define the context $C_{t}$ for any given response at time $t$ as all preceding conversational turns up to that point:
-$C_{t}=\{(\mathbf{x}_{i}^{(t)},\mathbf{y}_{i}^{(t-1)})|i<t\},$ (3)
+$C_{t}=\{(\mathbf{x}_{i}^{(t)},\mathbf{y}_{i}^{(t-1)})|i<t\},$
 where $\mathbf{x}_{i}^{(t)}$ denotes the sequence of student response till time $t$, and $\mathbf{y}_{i}^{(t-1)}$ denotes all tutor responses up to, but not including, the response at time $t$.
@@
 […]
+$\mathbf{y}_{t}^{\text{Rejected}}=\mathbf{y}_{t}^{\mathcal{A}_{S}},\text{and }f(\mathbf{y}_{t}^{\mathcal{A}_{T}})\neq f(\mathbf{y}_{t}^{\mathcal{A}_{S}}),$
+$\forall t\in\{1,\ldots,N\}\big{\}}$
 where $C_{t}=x,y_{w}=y_{choosen},y_{l}=y_{rejected}$ from DPO equation (1) from section 2.1, and the function $f$ is the key to distinguishing between chosen and rejected tutor responses.
@@
 This condition is captured through the function $f$—mapping responses to their pedagogical attributes:
-${f(\mathbf{y}_{i}^{\mathcal{A}})=\left(\mathrm{Eval}(\mathbf{y}_{i}^{\mathcal{A}}),\mathrm{Action}(\mathbf{y}_{i}^{\mathcal{A}}),\mathrm{Subprob}(\mathbf{y}_{i}^{\mathcal{A}})\right)}$ (5)
+${f(\mathbf{y}_{i}^{\mathcal{A}})=\left(\mathrm{Eval}(\mathbf{y}_{i}^{\mathcal{A}}),\mathrm{Action}(\mathbf{y}_{i}^{\mathcal{A}}),\mathrm{Subprob}(\mathbf{y}_{i}^{\mathcal{A}})\right)}$
   • $\text{Eval}(\mathbf{y}_{i}^{\mathcal{A}})$ extracts insights on the tutor’s grading of student’s solution or question understanding.
@@
 In this section, we provide a comprehensive overview of our dataset construction, experimental design, and the subsequent findings derived from the application of state-of-the-art LHP algorithms to train pedagogically-aligned LLMs.
+Model Metric SFT DPO IPO KTO
+Llama Acc 64 (62, 61, 70) 77 (74, 74, 84) 74 (71, 70, 81) 75 (73, 71, 82)
+F1 0.51 (0.54, 0.44, 0.55) 0.57 (0.62, 0.51, 0.59) 0.54 (0.57, 0.47, 0.57) 0.54 (0.58, 0.47, 0.58)
+Mistral Acc 65 (61, 61, 72) 74 (71, 69, 80) 71 (67, 67, 79) 72 (69, 69, 79)
+F1 0.47 (0.49, 0.41, 0.5) 0.52 (0.54, 0.47, 0.56) 0.5 (0.53, 0.45, 0.54) 0.51 (0.54, 0.46, 0.53)
+Zephyr Acc 23 (22, 21, 26) 73 (71, 70, 79) 72 (70, 69, 78) 75 (72, 72, 81)
+F1 0.26 (0.29, 0.23, 0.25) 0.55 (0.58, 0.45, 0.62) 0.52 (0.57, 0.45, 0.55) 0.56 (0.59, 0.47, 0.62)
+Table 2: SFT vs DPO, IPO, KTO.
+Each cell displays the average accuracy and F1 score across three classification fields: Evaluation of Student Response, Action Based on the Evaluation, and Subproblem State.
+We observe consistent improvements in both accuracy and F1 score for all three alignment algorithms—DPO, KTO, and IPO—compared to their SFT counterparts across the models Llama (Llama-3.1-8B-Instruct), Mistral (Mistral-7B-Instruct-v0.2), and Zephyr (zephyr-7b-beta).
+Notably, DPO and KTO consistently outperform IPO for all models.
 4.1 Dataset and Evaluation
@@
 Following the strategy outlined by CLASS, we set about creating these additional conversations.
-We used the CLASS scaffolding prompts to generate the problems, subproblems, and associated hints, which were based on biology topics from the OpenStax college books [18], consistent with CLASS’ context.
+We used the CLASS scaffolding prompt to generate the problems, subproblems, and associated hints, which were based on biology topics from OpenStax college textbooks (Clark et al., 2018).
 These problems seed the context of the conversations generated through GPT-4 in the next step using the CLASS conversational prompt.
@@
 Through this methodology, we successfully generated an additional $1128$ conversations, resulting in $1738$ conversations.
-Each conversation, on average, comprises approximately 8.2 rounds and 1311.0 words.
-This comprehensive dataset provides a solid platform for training and comparing SFT and LHP models.
-To create the training and testing set for SFT and LHP, we randomly sampled our conversational dataset into 3 partitions: the SFT training dataset $D_{c}$ (600 conversations), the LHP training dataset $D_{p}$ (600 conversations), and the test dataset $D_{t}$ (450 conversations).
-As each round within the conversations is processed into a separate data point, $D_{c}$ contains 4942 data points, $D_{p}$ contains 4921 data points, and $D_{t}$ contains 3701 data points.
+Each conversation, on average, comprises approximately $8$ rounds and $1311$ words.
+This dataset provides a solid platform for training and evaluating SFT and LHP models.
+To partition the training and testing set for SFT and LHP, we randomly sampled our conversational dataset into three partitions: SFT training dataset $D_{c}$ ($600$ conversations), LHP training dataset $D_{p}$ ($600$ conversations), and test dataset $D_{t}$ ($450$ conversations).
+As each round within the conversation is processed into a separate question-answer (QA) pair, $D_{c}$ contains $4942$ QA pairs, $D_{p}$ contains $4921$ QA pairs, and $D_{t}$ contains $3701$ QA pairs.
 For the evaluation, we focus on the three classification fields within the model JSON responses described in Section 3: Evaluation of Student Response, Action Based on Evaluation, and Subproblem State.
@@
 Hence, LLMs that achieve higher accuracy and F1 score in these fields are considered to provide more pedagogically aligned assitance to students.
-Model Metric SFT DPO IPO KTO
-Mistral Acc 35.0 (34,34,37) 74.0 (72,72,78) 57.7 (56,56,61) 75.3 (73,73,80)
-F1 0.34 (0.42,0.24,0.37) 0.56 (0.63,0.37,0.67) 0.48 (0.56,0.33,0.56) 0.55 (0.62,0.37,0.65)
-Zephyr Acc 10.7 (12,11,9) 74.0 (72,71,79) 21.7 (22,22,21) 74.7 (73,72,79)
-F1 0.14 (0.21,0.12,0.10) 0.56 (0.65,0.38,0.66) 0.26 (0.32,0.19,0.28) 0.54 (0.63,0.37,0.63)
-Table 2: SFT vs DPO, IPO, KTO.
-Each cell shows the average accuracy and average F1 score across three classification fields - Evaluation of student response, Action based on the evaluation, and Subproblem State, respectively.
-We observed that for both Mistral (Mistral-7B-Instruct-v0.2) and Zephyr (zephyr-7b-beta), all three alignment algorithms - DPO, KTO, and IPO - showed notable gains in accuracy and F1 score across all fields as compared to SFT.
-Across both models, DPO and KTO exhibited comparable performance, outperforming IPO.
 4.2 Models and and Training Details
-We experimented on two different LLMs: Mistral (Mistral-7B-Instruct-v0.2) and Zephyr (zephyr-7b-beta).
-We selected Mistral and Zephyr based on their top performance on diverse NLP benchmarks on OpenLLM leaderboard [19, 20] which evaluate scientific knowledge and education-based question answering capabilities: ARC (reasoning based science questions) [21], MMLU (multiple choice questions from 57 high school and college subjects across STEM, the humanities, the social sciences, and more) [22], and GSM8K (grade school math word problems) [23].
-These benchmarks offer a comprehensive evaluation on the models’ performance, especially in education setting.
+We experimented on three different LLMs: Llama (Llama-3.1-8B-Instruct), Mistral (Mistral-7B-Instruct-v0.2) and Zephyr (zephyr-7b-beta) with beta of 0.1 for DPO, IPO, and KTO across all models.
 Our experiments included training and evaluating SFT and LHP models.
-It took around 22 NVIDIA A6000 GPU hours for one cycle (SFT + Evaluation + LHP + Evaluation).
-A total of 17 cycles, which are 374 GPU hours, are needed to complete all the experiments.
-We conducted the SFT experiments using FastChat [24].
-For the choice of SFT hyperparameters, we refered to parameters used to instruct fine-tune Vicuna [24]: We employed a learning rate of 2e-5 with the AdamW optimizer [25], a batch size of 16, a cosine scheduler, a weight decay of 0.05, a warmup ratio of 0.1, and a total number of 3 epochs.
-We conducted the LHP experiments using TRL [26].
-For the choice of LHP hyperparamters, we referred to parameters used to align Zephyr [27]: We employed a learning rate of 1e-7 with the AdamW optimizer, a batch size of 16, a cosine scheduler, a weight decay of 0.05, a warmup ratio of 0.1, and a total number of 3 epochs.
+It took around $18$ NVIDIA A6000 GPU hours for one cycle (SFT + Data Generation + LHP + Evaluation).
+A total of $9$ cycles, which are $162$ GPU hours, are needed to complete all the experiments.
+We conducted the SFT experiments using FastChat (Zheng et al., 2023).
+For the choice of SFT hyperparameters, we refered to parameters used to instruct fine-tune Vicuna (Zheng et al., 2023): We employed a learning rate of $2\times 10^{-5}$ with the AdamW optimizer (Loshchilov and Hutter, 2019), a batch size of $16$, a cosine scheduler, a weight decay of $0.05$, a warmup ratio of $0.1$, and a total of 3 epochs.
+We conducted the LHP experiments using TRL (von Werra et al., 2020).
+For the choice of LHP hyperparamters, we use beta of $0.1$, a learning rate of $71e-$ 7 with the AdamW optimizer, a batch size of $16$, a cosine scheduler, a weight decay of $0.05$, a warmup ratio of $0.1$, and a total of $3$ epochs.
 4.3 Main Findings: SFT vs LHP
@@
 We measured the accuracy and F1 score across three classification fields in the responses: Evaluation of Student Response, Action based on Evaluation, and Subproblem State.
-We used the same LHP training setting as described in Section 4.2 with a beta of 0.1 for DPO, IPO, and KTO across all models.
-After applying the alignment algorithms, we observed significant performance improvements on Mistral and Zephyr SFT models.
-Specifically, Mistral demonstrated accuracy increase of 39%, 22.7%, and 39.7% from DPO, IPO, and KTO respectively, while Zephyr showed improvements of 63.3%, 11%, and 64% respectively (Table 2).
-We observed similar increase in F1 score as well: Mistral showed increase of 0.22, 0.14, and 0.21 from DPO, IPO, and KTO, and Zephyr showed improvements of 0.42, 0.12, and 0.4 respectively.
-Notably, for both Mistral and Zephyr, DPO and KTO significantly outperformed IPO.
-Our results are in alignment with the findings of similar studies [27], which have reported superior performance of DPO and KTO over IPO on Zephyr when the model was fine-tuned on preference datasets, such as Intel’s ORCA dataset pairs [28] and the Ultra-Feedback dataset [14].
-This experiment underscored the pivotal role of alignment algorithms in improving the accuracy and F1 score of SFT models.
-Mistral and Zephyr, the top performing models, exhibited pronounced improvements using LHP algorithms compared to SFT.
-4.4 Examining LHP through lens of perplexity
+We observed substantial performance improvements across all three SFT models—Llama, Mistral, and Zephyr.
+Specifically, Llama achieved accuracy gains of 13.1%, 10.1%, and 10.9% from DPO, IPO, and KTO, respectively.
+Mistral demonstrated increases of 8.7%, 6.1%, and 7.4%, while Zephyr showed improvements of 50.0%, 50.2%, and 52.0% (Table 2).
+A similar trend was observed in F1 scores: Llama exhibited increases of 6.3%, 2.7%, and 3.2%, Mistral showed gains of 5.6%, 3.7%, and 4.3%, and Zephyr saw improvements of 29.3%, 26.7%, and 30.2% for DPO, IPO, and KTO, respectively.
+Notably, DPO and KTO significantly outperformed IPO across all models.
+Also note these results are for a fixed beta of $0.1$.
+As we show in section 4.6, beta hyperparameter tuning can further boost the performance of LHP algorithms.
+This experiment shows the pivotal role of alignment algorithms in improving the accuracy and F1 score of SFT models.
+All models – Llama, Mistral and Zephyr exhibit pronounced improvements using LHP algorithms compared to SFT.
+4.4 Pedagogical Shifts: Perplexity Comparison of SFT and LHP
 For understanding the effectiveness of LHP algorithms compared to SFT, we propose a second preference dataset generation technique.
 This technique specifically allows us to perform a perplexity analysis to compare the LLM’s tendency to provide hints and guidance versus direct responses.
-The essence of pedagogical alignment is the LLM’s propensity to offer guidance (as seen in Actions 1 and 4) over direct solutions (as in Actions 2 and 5).
+Dataset Construction: The essence of pedagogical alignment is the LLM’s propensity to offer guidance (as seen in Actions 1 and 4) over direct solutions (as in Actions 2 and 5).
 By this definition, Action 2 can never precede Action 1, and likewise, Action 5 can never precede Action 4 (for more details about actions please refer to Table 1).
 With this critical insight, we use GPT-4 to generate pedagogically misaligned Actions (Action 2 for the first occurrence of Action 1 and Action 5 for the first occurrence of Action 4).
-This approach allows us to compare the perplexity of responses corresponding to Action 1 and its misaligned counterpart (Action 2), and similarly for Actions 4 and 5.
-Building upon the test split (450 conversations) of the originally generated dataset, we generate an alternative opposite action (Action 2 for Action 1, and Action 5 for Action 4) for the first occurrence of Action 1 or 4 in each conversation.
-We then use this newly constructed Action dataset to measure the perplexity of responses from the original instruct model (Base), its SFT fine-tuned counterpart (SFT), and its LHP version (DPO) (Table 3).
-Our analysis reveals that for both Mistral and Zephyr Base models, responses with Action 1 have higher perplexity than those with Action 2, and likewise for Actions 4 and 5.
-This suggests that the base models are more inclined to provide direct solutions than offer hints and guidance.
-The SFT process adjusts this behavior by increasing the perplexity of responses with Actions 2 and 5, thus making the models more likely to produce responses with Actions 1 and 4.
-However, SFT alone doesn’t fully optimize this behavior.
-As seen in the perplexity results for Mistral and Zephyr SFT models (Table 3), there is only a minimal decrease or even a slight increase in the perplexity of Action 1 and 4.
-This indicates that SFT, while useful, doesn’t significantly shift the model’s behavior towards providing more guidance over direct solutions.
-In contrast, LHP dramatically optimizes this behavior.
-For the Mistral model, the perplexity of Action 1 and 4 decreases from 2.05 and 2.14 to 1.89 and 1.96 (SFT vs DPO) respectively.
-Similarly, for the Zephyr model, the perplexity of Action 1 and 4 decreases from 2.09 and 2.18 (SFT) to 1.97 and 2.05 (DPO) respectively.
-This decrease in perplexity indicates a higher likelihood of the model providing guidance over direct solutions.
-Simultaneously, the perplexity of Action 2 and 5 increases and from 13.9 and 13.06 (Zephyr SFT) to 14.9 and 14.38 (Zephyr DPO).
-This increase in perplexity indicates a lower likelihood of the model providing direct solutions, further emphasizing the model’s shift towards providing guidance.
+This approach allows us to compare the perplexity of responses corresponding to Action 1 and its misaligned counterpart Action 2, and similarly for Actions 4 and 5.
+We then use this newly constructed Action dataset to evaluate the perplexity of responses generated by the original instruct model (Base) to its SFT and LHP variants.
+Figure 3: Comparison of multi-round performance for SFT vs.
+LHP methods (DPO, IPO, KTO) across Llama, Mistral, and Zephyr models.
+The graphs illustrate average accuracy over 8 conversation rounds, revealing the superior performance of LHP methods in maintaining pedagogical alignment across extended conversation context.
+Figure 4: Performance of LHP algorithms (DPO, IPO, and KTO) as a function of beta. Our results indicate that KTO outperforms both DPO and IPO with optimal beta hyperparameter search.
+Perplexity Findings: Our analysis, as shown in table 3), reveals interesting patterns across the three models: (A) Llama: The base model shows similar perplexity across all actions, indicating no strong preference for guidance or direct solutions.
+SFT increases perplexity for Actions 2 and 5, shifting towards more guidance.
+Both DPO and KTO further optimize this behavior, with DPO showing the lowest perplexity for Actions 1 and 4; (B) Mistral: The base model shows slightly higher perplexity for Actions 1 and 2.
+SFT dramatically increases perplexity for Actions 2 and 5, strongly favoring guidance.
+DPO and KTO further reduce perplexity for Actions 1 and 4, with DPO showing slightly better optimization; (C) Zephyr: The base model shows relatively uniform perplexity across actions.
+SFT significantly increases perplexity for Actions 2 and 5.
+DPO and KTO further optimize this behavior, with DPO showing lower perplexity for Actions 1 and 4, while KTO shows higher perplexity for Actions 2 and 5.
+These results demonstrate that SFT partially shifts model behavior towards providing more guidance, but LHP methods (DPO and KTO) further optimize this behavior.
+For Mistral and Zephyr, LHP methods dramatically increase the perplexity gap between guidance actions (1 and 4) and direct solution actions (2 and 5), indicating a significant shift towards pedagogical alignment.
+The perplexity analysis provides compelling evidence for the effectiveness of LHP methods in achieving pedagogical alignment, consistently outperforming SFT across all three models.
 Model A1 A2 A4 A5
-Mistral (Base) 2.11 2.06 2.00 1.97
-Mistral (SFT) 2.05 5.55 2.14 5.40
-Mistral (DPO) 1.89 5.48 1.96 5.34
-Zephyr (Base) 2.05 2.09 2.03 2.07
+Llama (Base) 2.02 2.04 2.07 2.01
+Llama (SFT) 1.77 2.44 1.87 2.45
+Llama (DPO) 1.70 2.45 1.78 2.43
+Llama (KTO) 1.74 2.48 1.81 2.46
+Mistral (Base) 2.11 2.06 2.01 1.97
+Mistral (SFT) 2.06 5.51 2.14 5.37
+Mistral (DPO) 1.86 5.49 1.93 5.31
+Mistral (KTO) 1.90 5.01 1.98 4.93
+Zephyr (Base) 2.05 2.09 2.04 2.07
 Zephyr (SFT) 2.09 13.9 2.18 13.06
-Zephyr (DPO) 1.97 14.9 2.05 14.38
-Table 3: Perplexity of responses.
-Perplexity analysis reveals that the base models of Mistral and Zephyr are more inclined to provide direct solutions (Action 2 and 5) than guidance and hints (Action 1 and 4).
-The SFT process partially shifts this behavior, increasing the perplexity on responses with Action 2 and 5, thus making the models more likely to produce responses with Action 1 and 4.
-However, the SFT process alone doesn’t fully optimize this behavior, as seen by the minimal decrease or slight increase in the perplexity of Action 1 and 4.
-In contrast, DPO dramatically optimizes this behavior by reducing perplexity on Action 1 and 4 while simultaneously increasing perplexity on Action 2 and 5.
-This shift in perplexity indicates a significant behavioral shift in the LLM, making it more aligned with the role of a tutorbot by favoring guidance over direct solutions.
+Zephyr (DPO) 1.91 16.44 1.97 15.66
+Zephyr (KTO) 1.97 16.76 2.03 16.16
+Table 3: Perplexity analysis of model responses corresponding to different actions.
+Lower perplexity for Actions 1 and 4 indicates a higher likelihood of providing scaffolded guidance like hints, while higher perplexity for Actions 2 and 5 suggests a lower tendency to offer direct solutions.
+Results show SFT partially improves pedagogical alignment, while LHP methods (DPO, KTO) further optimize it by widening the perplexity gap between guidance (Actions 1, 4) and direct solution (Actions 2, 5) responses.
 For more details about actions please refer to Table 1.
-4.5 Effect of Beta on DPO and IPO
-Figure 3: DPO and IPO performance as a function of beta.
-We found that classification accuracy and F1 score using both DPO and IPO were significantly influenced by beta hyper-parameter adjustments.
-Specifically, a beta of 0.3 optimized results for DPO, with Mistral and Zephyr achieving around 75% accuracy and 0.56 F1 score, while a higher beta of 0.9 notably reduced accuracy and F1 score.
-In contrast, IPO performed best with a beta of 0.1 and increasing beta decreased accuracy and F1 score.
-In this experiment, we analyzed the impact of the beta parameter on the performance of DPO and IPO alignment algorithms.
+4.5 Pedagogical Consistency Over Time: SFT vs. LHP in Extended Conversations
+To assess the robustness and consistency of pedagogical alignment across extended conversations, we conducted an in-depth analysis of model accuracy across multiple conversation rounds.
+This analysis is important because as conversations progress, the context becomes more complex and lengthy, potentially challenging the model’s ability to maintain consistent pedagogical alignment.
+The analysis reveals striking differences in pedagogical alignment between SFT and LHP methods across all three models.
+As can be seen in Figure 3, Llama’s SFT model shows a clear decline in alignment as conversations progress, dropping from $75\%$ to $60\%$ by round $8$, while LHP methods maintain consistent alignment around $75\%$.
+Mistral and Zephyr demonstrate even more pronounced contrasts.
+Their SFT models struggle with pedagogical alignment in extended conversations, with Mistral fluctuating between $63-69\%$ and Zephyr performing poorly at $11-30\%$.
+In stark contrast, all LHP approaches (DPO, IPO, KTO) for both models maintain high pedagogical alignment above $70\%$ throughout, with Zephyr showing the most dramatic improvement over its SFT counterpart.
+This consistent performance of LHP methods across all models, particularly in later conversation rounds, underscores their effectiveness in maintaining pedagogical alignment even as context complexity increases.
+These results demonstrate that LHP-aligned models maintain high pedagogical alignment even with increasing context length, unlike SFT models which tend to degrade in performance over extended conversations.
+This robustness is crucial for effective tutoring, as it ensures the model can provide reliable guidance throughout a problem-solving session, regardless of its context length.
+4.6 Effect of Beta on LHP Algorithms
+In this experiment, we analyze the impact of the beta parameter on the performance of LHP algorithms.
 Beta is a key hyperparameter in all three alignment algorithms (DPO, IPO, and KTO).
-It controls the strength of the Kullback-Leibler (KL) divergence penalty between the trained model and the reference model within the loss function.
-Nevertheless, the choice of beta depends on the specific model and dataset and thus necessitates an empirical study.
-Similar to experiments in Section 4.3, we measured the average accuracy and F1 score across the three classification fields.
-We found that the performance of DPO and IPO is sensitive to the choice of beta.
-As depicted in Figure 3, for DPO, we observed that a beta value of 0.3 yielded the best results: Mistral and Zephyr achieved an average accuracy of 74.7% and 75% and an average F1 score of 0.55 and 0.56.
-Conversely, a larger beta value, such as 0.9, diminished the performance of the models.
-Zephyr suffered a large loss of 25.7 % in average accuracy from using 0.9 instead of 0.3, and a loss of 0.13 in average F1 score.
-As opposed to IPO, we observed that a beta of 0.1 produced the best performance (Figure 3).
-A larger choice of beta in IPO almost always came with a loss of performance.
-As beta increased, Mistral’s accuracy dropped from 57.7% to 36 %, and its F1 score decreased from 0.55 to 0.49, while Zephyr’s accuracy dropped from 21.7 % to 11.3 %, and its F1 score decreased from 0.56 to 0.43.
-Our results corroborate the evidence from related studies [27], demonstrating that various alignment algorithms exhibit distinct sensitivities to beta values, necessitating tailored beta selections.
+It controls the strength of the Kullback-Leibler divergence penalty between the trained model and the reference model within the loss function.
+The choice of beta depends on the specific model and dataset and thus necessitates an empirical study.
+As we show next, beta hyperparameter tuning can significantly affect the performance of LHP algorithms, with KTO emerging as the top performer.
+Similar to experiments in Section 4.3, we measure the average accuracy and F1 score across the three classification fields.
+We find that the performance of all three algorithms is sensitive to the choice of beta, but to varying degrees and with different optimal values for each algorithm and model.
+For DPO, we observe that lower beta values (0.1 or 0.3) generally yield the best results.
+Llama achieved the highest average accuracy of 77.3% and average F1 score of 0.57 with beta as 0.1.
+Mistral performed best with beta = 0.1, reaching 74.1% accuracy and 0.52 F1 score.
+Zephyr showed similar performance for beta = 0.1 and 0.3, with a slight edge for 0.1 (73.4% accuracy, 0.55 F1 score).
+IPO demonstrated less sensitivity to beta changes, especially for Llama and Mistral.
+Llama’s performance peaked at beta = 0.6 with 74.9% accuracy and 0.54 F1 score.
+Mistral showed minimal variation, with slightly better results at beta = 0.6 (71.4% accuracy, 0.51 F1 score).
+Zephyr performed best with beta = 0.3, achieving 72.6% accuracy and 0.55 F1 score.
+KTO emerged as the top performer, showing consistent improvement with increasing beta values across all models.
+For Llama, performance peaked at beta = 0.9 with 78.1% accuracy and 0.57 F1 score.
+Mistral also showed best results at beta = 0.9, reaching 74.7% accuracy and 0.51 F1 score.
+Zephyr demonstrated steady improvement, with optimal performance at beta = 0.9 (75.7% accuracy, 0.55 F1 score).
+Notably, with optimal beta values, KTO outperforms both DPO and IPO across all models, emphasizing its effectiveness in pedagogical alignment tasks when properly tuned.
+Our findings underscore the necessity of careful hyperparameter tuning in LHP algorithms to achieve optimal pedagogical alignment.
+They also suggest that KTO, when properly tuned, may be particularly well-suited for tasks requiring sustained pedagogical alignment.
 5 Conclusion
-In this paper, we have investigated the application of learning from human preferences (LHP) methods to align LLMs with educational objectives, aiming to foster optimal student learning.
+In this paper, we have investigated the application of learning from human preferences (LHP) methods to align LLMs with educational goals, aiming to foster optimal student learning outcomes.
 By constructing a preference dataset specifically designed for pedagogical alignment, we have laid the foundation for enhancing the effectiveness of LLMs in educational contexts.
-Our experimental results, derived from applying LHP alignment algorithms on state-of-the-art open-source LLMs, demonstrate the superiority of LHP methods over the traditional supervised fine-tuning (SFT) approach, significantly improving the alignment of LLMs with pedagogical objectives.
+Our experimental results, derived from applying LHP alignment algorithms on state-of-the-art open-source LLMs, demonstrate the superiority of LHP methods over standard supervised fine-tuning (SFT), significantly improving the alignment of LLMs with pedagogical objectives.
 Our another key contribution is the development of a novel approach for evaluating the pedagogical effectiveness of LLMs.
-By curating a second preference dataset that compares the perplexity of responses offering scaffolding versus those providing direct solutions, we have introduced a new methodology that quantifies the extent to which LLMs prioritize pedagogically effective actions.
+By curating a preference dataset that compares the perplexity of responses offering scaffolding versus those providing direct solutions, we have introduced a new methodology that quantifies the extent to which LLMs prioritize pedagogically effective actions.
 This approach opens up new possibilities for assessing and optimizing the performance of LLMs in educational settings, providing a valuable tool for researchers and practitioners alike.
+6 Limitations
+While our methodology shows promising results, there are certain inherent limitations that are challenging to overcome.
+One such limitation is the dynamic and complex nature of the educational landscape.
+The effectiveness of our approach may vary depending on a multitude of factors, including the diversity of student learning styles, the evolving nature of educational content, and the rapid advancements in AI technology.
+Additionally, computational constraints may also limit the scalability of our approach in larger, more diverse educational settings.
+More comprehensive user studies can assess the impact of these aligned models in real-world educational settings.
+7 Ethics and Risks
+The use of LLMs in education comes with ethical considerations and potential risks.
+These include the need to ensure the models do not amplify biases, the importance of protecting sensitive student data, and the necessity of maintaining human guidance in the learning process.
+In addition, the impact of LLMs on learning outcomes should be continuously evaluated to ensure their effectiveness and avoid unintended negative consequences.
+Acknowledgments
+This work was supported by NSF grant 1842378, ONR grant N0014-20-1-2534, AFOSR grant FA9550-22-1-0060, a Vannevar Bush Faculty Fellowship, and ONR grant N00014-18-1-2047.